### v1

# Modality Collapse as Mismatched Decoding: Information-Theoretic Limits of Multimodal LLMs

Abstract
[…]
We formalize this as a mismatched decoder problem: a decoder trained on text can only extract information along text-aligned directions. Accessible information is bounded by the Generalized Mutual Information (GMI), with degradation scaling with distributional distance and decoder sensitivity. The bound is a property of the decoder’s scoring rule, not of any particular architecture; it applies whether non-text inputs arrive through a learned projection, a discrete codebook, or no explicit adapter at all. We validate this across five models spanning speech and vision. A controlled experiment (two Prismatic VLMs differing only in encoder text-alignment) confirms the bottleneck is the decoder’s scoring rule, not the encoder or projection. A LoRA intervention demonstrates the fix: training with an emotion objective improves emotion accessibility (++7.5%) without affecting other attributes, confirming that the training objective determines what becomes accessible.
1 Introduction
Multimodal LLMs follow a common recipe: an encoder processes the non-text input (speech, image), a learned projection maps it into the LLM’s embedding space, and the LLM generates text (Liu et al., 2023; Dai et al., 2023; Fixie AI, 2024; Chu et al., 2024; Karamcheti et al., 2024). By standard benchmarks these systems appear to work well: speech models transcribe accurately, vision models caption faithfully. But this success is a mirage. These systems excel at tasks requiring only text-correlated information (the words spoken, the objects visible) while silently discarding everything else. A speech model that transcribes perfectly may have no access to who is speaking or how they feel; a vision model that names every object may be blind to spatial layout or texture. This paper shows that the loss is structural and predictable.
[…]
Worse, modality-specific information degrades decoding. Removing 64% of the variance in modality-specific directions improves cross-entropy loss; the decoder works better with less information. Lexical probe accuracy nearly doubles from the adapter to the final LLM layer (the decoder amplifies what it knows how to use), while speaker probe accuracy drops by up to 39% over the same path. The mechanism is not active suppression: the decoder simply has no learned response to non-text directions. It is indifferent to them, but that indifference has consequences, because the decoder treats unfamiliar variance as noise. Where the collapse manifests depends on the encoder: non-aligned encoders (Whisper, DINOv2) pass rich representations the decoder is blind to; text-aligned encoders (CLIP, SigLIP) discard non-textual information before it reaches the decoder. Either way, the information is lost (Appendix Table 6).
Our thesis. We formalize this as a mismatched decoder problem (Merhav et al., 1994; Scarlett et al., 2020). The LLM’s scoring rule was learned from text, so it can only extract information along text-aligned directions. Accessible information is bounded by the Generalized Mutual Information (GMI), not the standard mutual information. We call the difference the information accessibility gap. The gap is not about architecture; it is about the scoring rule. No explicit adapter module is required: any system where a text-trained decoder processes non-text representations faces the same constraint, whether the representations arrive through a learned projection, a discrete codebook, or the LLM’s own layers. At inference, every trained model has fixed weights: the scoring rule is whatever the training procedure produced. A model trained predominantly on text has a text-shaped scoring rule regardless of whether the LLM was frozen or fine-tuned during training. What matters is the training objective, not which components were updated (§6.1).
What determines the size of the gap? A linear probe simply measures whether the information it was trained to detect is present. The decoder is a different instrument entirely. It is a deep transformer trained on text, and it is highly sensitive to its input, responding to variance in every direction, including non-text directions (we confirm this empirically via gradient isotropy in §5.5). But its response to non-text variance is unproductive: the decoder interprets unfamiliar structure as noise, and that noise disrupts its text processing. This is why removing modality-specific directions helps: not because the decoder ignores them, but because it responds to them in ways that degrade output. We formalize this: degradation scales with both the distributional distance between modal and text representations and the decoder’s sensitivity to that distance (Theorem 2), and the decoder’s sensitivity is up to $30\times$ that of a probe (Theorem 3, §A.4). Text-aligned encoders reduce the distance (§5.3), but only by discarding non-textual information upstream. The collapse happens earlier, not less.
Contributions. (1) We formalize modality collapse as mismatched decoding and prove that accessible information is bounded by the GMI, with degradation governed by distributional distance and decoder sensitivity (§4). (2) Across five models and two modalities, we demonstrate the information accessibility gap: non-text information survives through the LLM but degrades decoding because the decoder has no learned use for it (§5.2, §5.4). (3) A controlled experiment with Prismatic VLMs (same architecture, same LLM, only encoder text-alignment differs) isolates the scoring rule as the causal variable (§5.3). (4) A LoRA intervention confirms the prescription: training with an emotion objective teaches the decoder to respond to that dimension (++7.5%), while a text-centric objective leaves it indifferent (§6.1).
2 Related work
Multimodal LLMs and the adapter paradigm. The strategy of projecting encoded features into an LLM’s embedding space has become standard for both vision (Liu et al., 2023; Dai et al., 2023; Zhu et al., 2024; Karamcheti et al., 2024; Tong et al., 2024) and speech (Fixie AI, 2024; Chu et al., 2024). These systems achieve strong performance on text-centric benchmarks but struggle with tasks requiring modality-specific information, an observation that has been noted empirically but not explained theoretically.
Modality gap. Liang et al. (2022) showed that CLIP embeddings of different modalities occupy distinct regions of the shared space, separated by a persistent “modality gap.” Huh et al. (2024) observed convergence of representations across modalities. Our work complements these geometric observations with an information-theoretic mechanism: the gap matters not because information is destroyed, but because the decoder cannot access information in unfamiliar directions.
Mismatched decoding. The theory of mismatched decoders (Merhav et al., 1994; Lapidoth, 1996; Scarlett et al., 2020) characterizes the information rate achievable when a decoder is designed for a different channel than the one actually used. We import this framework into the multimodal LLM setting: at inference, any LLM whose scoring rule was shaped by text-dominated training is a mismatched decoder for non-text inputs. The mismatch is determined by the training objective, not by which components were updated.
Representation probing. Linear probes are a standard tool for measuring information content in neural representations (Alain & Bengio, 2017; Conneau et al., 2018; Hewitt & Manning, 2019; Belinkov, 2022). We use probes not to claim information is “encoded” in a representation, but to demonstrate the gap between information present (probe-recoverable) and information accessible (decoder-extractable).
Rate-distortion in ML. Blau & Michaeli (2019) introduced the rate-distortion-perception tradeoff, showing that perceptual quality constraints can increase the rate needed for lossy compression. Our theoretical framework shares the insight that the output distribution shape (here, conformity to the text manifold) imposes additional costs beyond standard rate-distortion.
3 Problem formulation
Having described the phenomenon, we now formalize the architecture and the mismatch that causes it.
Architecture. We consider the standard multimodal LLM pipeline:
$X\xrightarrow{E_{\phi}}H\xrightarrow{P_{\theta}}Z\xrightarrow{L_{\psi}(\cdot,C)}\hat{Y}$ (1)
where $X$ is a non-text input (speech waveform or image), $E_{\phi}$ is a pre-trained encoder (e.g., Whisper (Radford et al., 2023), CLIP (Radford et al., 2021)), $P_{\theta}$ is a learned adapter (linear or MLP projection), $Z\in\mathbb{R}^{d}$ are the hidden states consumed by the LLM $L_{\psi}$, $C$ is the text context (prompt), and $\hat{Y}$ is the generated output. At inference, all components are fixed: the scoring rule $q_{\psi}$ is whatever the training procedure produced. In the common adapter paradigm, only $P_{\theta}$ is trained while $E_{\phi}$ and $L_{\psi}$ are pre-trained; in end-to-end approaches, all components are jointly trained. Either way, at inference the decoder’s scoring rule is fixed, and our theory describes what that fixed scoring rule can extract. The specific form of $P_{\theta}$ varies (linear, MLP, Q-Former (Dai et al., 2023), perceiver, codebook), but the mismatch depends on the scoring rule, not the adapter type (§6.1).
Text law and modal law. Let $q_{\psi}(y|z,c)=L_{\psi}(z,c)$ denote the LLM’s next-token distribution at inference. During pre-training, the LLM processes hidden states drawn from the text law $P_{\!T}(C,Z,Y)$. When processing adapted non-text inputs, the hidden states follow a different modal law $P_{\!M}(C,Z,Y)$.
Assumption 1 (Shared marginal).
The text and modal laws share the same prompts and target tokens: $P_{\!M}(C,Y)=P_{\!T}(C,Y)$.
In other words, the task is the same (e.g., “transcribe this speech”), and only the internal representations $Z$ differ. This isolates the effect of the representation distribution from confounds in the task itself.
Modality collapse. We define modality collapse as the systematic degradation or stagnation of non-text information through a text-trained decoder, depending on the mismatch intensity. For an attribute $A_{\tau}$ of type $\tau$ (e.g., speaker identity, emotion), the information accessibility gap is:
$\Delta_{\text{access}}(\tau)=I(Z;A_{\tau})-\mathrm{GMI}_{P_{\!M}}(A_{\tau}\mid\psi)$ (2)
where $I(Z;A_{\tau})$ is the mutual information (recoverable by an optimal probe) and $\mathrm{GMI}_{P_{\!M}}(A_{\tau}|\psi)$ is the information actually extractable by the decoder’s fixed scoring rule. When $\Delta_{\text{access}}=0$, the decoder extracts everything the probe can; when $\Delta_{\text{access}}$ is large, information is present but inaccessible.
Text-aligned encoders as correlation mappings. Contrastive encoders like CLIP (Radford et al., 2021) and SigLIP (Zhai et al., 2023) do not merge the modality manifolds (the modality gap persists (Liang et al., 2022)), but they align the informative directions: attributes that co-occur with text descriptions occupy dimensions also used by text representations. When such an encoder feeds an LLM, the modal law $P_{\!M}$ has high overlap with $P_{\!T}$ along informative dimensions, so the distributional distance is small and the accessibility gap is correspondingly small (§A.2 formalizes this via the Wasserstein distance). Non-contrastive encoders like DINOv2 (Oquab et al., 2024) lack this alignment: their informative directions may be orthogonal to $P_{\!T}$, making the information effectively inaccessible to the decoder. We return to this in §5.3, where the Prismatic comparison provides direct evidence.
4 Theoretical framework
The formulation above makes the information loss quantifiable. We establish three results (formal statements, assumptions, and proofs in Appendix A):
1. The decoder has a ceiling. Standard mutual information $I(Z;Y)$ measures how much information $Z$ contains about $Y$ assuming an optimal decoder. But the LLM is not optimal for non-text inputs. The relevant quantity is instead the Generalized Mutual Information (GMI) (Merhav et al., 1994; Scarlett et al., 2020), which measures the maximum rate extractable by a fixed decoder with a given scoring rule. Regardless of how much information a probe can extract from $Z$, the decoder can never exceed $\mathrm{GMI}_{P_{\!M}}(q_{\psi})$ (Theorem 1).
2. The ceiling drops with distributional distance. When modal representations $P_{\!M}$ differ from the text representations $P_{\!T}$ the decoder was trained on, the GMI degrades. We bound this degradation (Theorem 2): it scales with the product of two empirically measurable quantities: (i) the Wasserstein distance $W_{\!1}(P_{\!M},P_{\!T})$, measuring how far modal representations are from text representations, and (ii) the decoder’s Lipschitz constant $L_{\log}$, measuring how sensitive its log-probability is to shifts in the input. The bound makes a clear prediction: models with larger $L_{\log}\cdot W_{\!1}$ should show more degradation, which we validate in §5.5.
[…]
5 Experimental validation
The bound in Theorem 2 makes three quantities empirically measurable: (i) the Lipschitz constant $L_{\log}$, estimated via gradient norms, measuring the decoder’s sensitivity to input perturbations; (ii) the Wasserstein distance $W_{\!1}(P_{\!M},P_{\!T})$, estimated between modal and text hidden states at matched layers; and (iii) the mode alignment spectrum $\tilde{\alpha}(u_{k})$, which decomposes $W_{\!1}$ geometrically by identifying which directions carry the mismatch. We validate each in turn.
5.1 Setup
Models. We study five multimodal LLMs spanning two modalities (Table 1). For speech: Ultravox-v0.6 (Fixie AI, 2024) (Whisper encoder, MLP adapter with SwiGLU, Llama-3.1-8B) and Qwen2-Audio-7B (Chu et al., 2024) (Whisper encoder, linear adapter, Qwen2-7B). For vision: LLaVA-v1.5-7B (Liu et al., 2023) (CLIP-ViT, MLP adapter, Vicuna-7B) and two Prismatic VLMs (Karamcheti et al., 2024) sharing the same MLP adapter and Vicuna-7B backbone but differing only in their vision encoder: DINOv2 (not text-aligned) and SigLIP (text-aligned via contrastive loss).
Model Encoder Text-aligned? LLM
Ultravox Whisper No Llama-3.1-8B
Qwen2-Audio Whisper No Qwen2-7B
LLaVA CLIP-ViT Yes Vicuna-7B
Prismatic-D DINOv2 No Vicuna-7B
Prismatic-S SigLIP Yes Vicuna-7B
Table 1: Models studied. Prismatic-D and Prismatic-S share the same architecture, adapter, training recipe, and LLM backbone; only the vision encoder differs.
Datasets. Speech: LibriSpeech (Panayotov et al., 2015) (2,620 samples, 50 words, 40 speakers), CREMA-D (Cao et al., 2014) (7,442 samples, 6 emotions, 91 speakers), ESC-50 (Piczak, 2015) (2,000 samples, 50 sound classes). Vision: MS-COCO (Lin et al., 2014) (5,000 images, 69 object categories, 12 super-categories).
Probing protocol. We extract representations at four hook points per model: encoder output, adapter output, LLM layer 16, and LLM final layer (after LayerNorm). Representations are mean-pooled across sequence positions (restricting to audio-only tokens yields nearly identical results; Appendix Table 9). We train $\ell_{2}$-regularized logistic regression probes (5 seeds, 80/20 stratified split, $z$-score normalization) for each combination of hook point, information type, and dataset.
Information types. Speech: lexical (word identity), speaker (speaker identity), emotion (6-way), acoustic (sound class). Vision: lexical (caption-based word prediction), object category (69-way), super-category (12-way).
The Prismatic pair (shared architecture, different encoder) and the two speech models (shared encoder, different LLM) provide controlled contrasts that isolate specific variables.
5.2 Modality-specific structure interferes with decoding
We directly test whether the decoder’s scoring rule functionally depends on modality-specific (MS) directions by projecting them out at inference time and measuring the change in cross-entropy. Unlike probing, which tests representational presence, this is a direct intervention on the decoder’s input.
We decompose the adapter output into eigenmodes (principal directions of variance) and ask: how much of this variance does the text distribution also use? Each eigenmode $u_{k}$ gets a text alignment score $\tilde{\alpha}(u_{k})=u_{k}^{\top}\Sigma_{T}u_{k}/\lambda_{k}$, the fraction of its variance that text representations share (§5.6). Modes with $\tilde{\alpha}<0.5$ are modality-specific (MS): most of their variance comes from the non-text modality. We then surgically remove selected directions from the decoder’s input ($z^{\prime}=z-\sum_{k\in S}(z^{\top}u_{k})\,u_{k}$) and measure the effect on cross-entropy. We compare three conditions (200 samples per model): (i) ablate all MS modes, (ii) ablate a matched number of top text-aligned (TA) modes, (iii) no ablation.
[…]
Non-aligned encoders
Ultravox MS (all) 11 63.6 -1.4
TA-matched 11 29.4 -0.3
Random 11 2.8 ++0.3
Prismatic-D MS (all) 53 71.0 -11.1
TA-matched 47 29.0 -0.07
Random 53 39.1 -5.1
Text-aligned encoder
Prismatic-S MS (all) 14 24.6 -0.5
TA-matched 14 36.2 -0.7
Random 14 10.6 -0.2
Table 2: Causal ablation at the adapter output. Eigenmodes are classified as modality-specific (MS; $\tilde{\alpha}<0.5$) or text-aligned (TA), and matched subsets are projected out. Random controls ablate the same number of modes chosen uniformly at random (seed-averaged). For non-aligned encoders, removing MS modes improves loss while removing equal-sized TA subsets has negligible effect; the contrast is $160\times$ for Prismatic-D (t=25.2t=-25.2 vs. t=0.4t=-0.4). Random ablation gives intermediate effects proportional to variance removed. For SigLIP, few modes are MS and all effects are $1\%$.
MS structure is not merely unused; its presence degrades decoding. Unlike concept erasure methods such as LEACE (Belrose et al., 2023), which project out directions predictive of a labeled attribute, our ablation targets directions identified by distributional mismatch between modal and text covariances, without requiring any concept labels. The question is not whether a concept can be erased, but whether the decoder uses these distribution-mismatched directions at all.
Table 2 reveals a stark asymmetry across all three models. For Ultravox, removing 11 MS modes (63.6% of variance) slightly improves cross-entropy (-1.4%, t=7.0t=-7.0), while removing a matched TA subset has negligible effect (-0.3%, t=2.2t=-2.2). The Prismatic-D result is even more striking: ablating 53 MS modes (71% of variance) reduces loss by 11.1% (t=25.2t=-25.2), while ablating the top 47 TA modes (29% of variance) has no measurable effect (-0.07%, t=0.4t=-0.4, $p=0.69$). The decoder is functionally indifferent to the directions where most adapter variance concentrates, precisely as the mismatched-decoder theory predicts.
For Prismatic-S (text-aligned), only 14 of 100 modes are MS (24.6% of variance) and all ablation effects are below 1%, consistent with SigLIP producing representations the decoder already expects.
5.3 Controlled experiment: text-aligned vs. non-aligned encoders
The causal ablation establishes that MS directions interfere with decoding. We now isolate why some models have more MS variance than others, using the Prismatic pair as a controlled comparison that holds everything constant except encoder text alignment.
Both Prismatic VLMs share the same MLP adapter, Vicuna-7B backbone, and training recipe. The only difference is the vision encoder: DINOv2 (Oquab et al., 2024) (no text alignment) versus SigLIP (Zhai et al., 2023) (contrastive text–image loss). If our theory is correct, SigLIP should produce representations closer to $P_{\!T}$, yielding smaller $W_{\!1}(P_{\!M},P_{\!T})$ and less information loss.
The results confirm the prediction (Appendix Figure 1, Table 10). With DINOv2, visual information passes through the LLM essentially unchanged (-0.0% object category). With SigLIP, the same information types improve (++1.9% object category, ++2.3% super-category). This pattern replicates across modalities: in speech, Whisper-based encoders show dramatic lexical recovery (++92–95%) alongside speaker collapse (-8% to -39%); in LLaVA (CLIP encoder, text-aligned), all information types improve (Appendix Table 7).
5.4 The information accessibility gap in practice
The causal ablation (§5.2) establishes that the decoder does not use MS directions, and the controlled comparison (§5.3) traces this to encoder text alignment. Probes confirm the information has not been destroyed (Appendix Tables 75).
[…]
Three non-textual visual attributes (object count, size, spatial spread) derived from COCO annotations but absent from captions confirm the pattern in a milder mismatch regime ($L_{\log}\cdot W_{\!1}\approx 13$$54$ vs. $162$ for speech), explaining stagnation rather than degradation (Appendix Table 10).
5.5 Validating the bound
Theorem 2 predicts that GMI degradation scales with $L_{\log}\cdot W_{\!1}$. We estimate both quantities empirically (full results in Appendix Table 8). $L_{\log}$ is finite for all nine model-dataset combinations, confirming the Lipschitz log-score assumption (Assumption 2, Appendix A.2). Three patterns emerge: (i) $L_{\log}$ is content-dependent, decreasing as audio moves away from text ($9.08$ speech $\to$ $3.66$ environmental sounds in Ultravox); (ii) $L_{\log}$ varies $30\times$ across architectures ($0.29$ LLaVA to $9.08$ Ultravox); (iii) Prismatic-S has $2.8\times$ higher $L_{\log}$ than Prismatic-D ($1.12$ vs. $0.40$), consistent with the LLM being more responsive to text-aligned representations. $W_{\!1}$ ranges from 17.8 to 50.4 across models.
Gradient isotropy. The mean gradient norm along modality-specific directions ($g_{\text{MS}}$) is nearly identical to that along text-aligned directions ($g_{\text{TA}}$): $g_{\text{TA}}/g_{\text{MS}}=0.94\times$, Spearman $r=0.11$, $p=0.28$. This validates the scalar $L_{\log}$ assumption: the decoder responds with similar magnitude to perturbations in any direction. But MS modes carry structured variance that pushes predictions away from text distributions, while TA modes carry variance the decoder was trained to exploit. Same sensitivity, different semantic effect.
The product $L_{\log}\cdot W_{\!1}$ is largest for Ultravox ($162$), which shows the most dramatic degradation (-39% speaker), and smallest for LLaVA ($13.4$), where all information types improve. Prismatic-S has a higher product ($53.7$) than Prismatic-D ($20.2$) despite showing less degradation; this reflects higher decoder responsiveness to text-aligned representations rather than greater mismatch. The bound’s prefactor is also smaller for Prismatic-S because its effective support diameter is smaller ($D=16.2$ vs. $35.7$).
The raw bound uses the ambient representation diameter $D$, which makes the prefactor vacuously large. The support-restricted bound (Appendix A.3) allows replacing $D$ with the effective support diameter $D_{\text{eff}}$, estimated via the participation ratio of the pooled representation covariance. This yields $L_{\log}\cdot D_{\text{eff}}\in[3,7]$, making the prefactor moderate ($\sim$20–1100). The bound is structural rather than tight: it identifies $L_{\log}\cdot W_{\!1}$ as the governing quantity and correctly predicts which models collapse most.
Per-type prediction. The theory correctly predicts the aggregate pattern: text-aligned information recovers or improves, while non-text types degrade or stagnate. A finer prediction (degradation correlating with each type’s MS-variance share) does not hold at this granularity (Spearman $\rho=-0.40$, $p=0.60$), likely because the content-dependent $L_{\log}$ (which varies $2.5\times$ across audio types) modulates degradation beyond geometric alignment alone.
5.6 Mode alignment
The bound explains how much information is lost; mode alignment explains which directions carry the mismatch. We decompose the adapter output into its principal directions of variance (eigenmodes of $\Sigma_{M}$) and measure, for each direction, the fraction of its variance that the text distribution shares: $\tilde{\alpha}(u_{k})=u_{k}^{\top}\Sigma_{T}u_{k}/\lambda_{k}$. A score near 1 means the direction is shared with text; near 0 means it is unique to the modality (Appendix Table 11, Figure 2).
For non-aligned encoders, the largest direction (Mode 0) is modality-specific ($\tilde{\alpha}\leq 0.034$), capturing 51–79% of adapter variance. Most of what the adapter transmits is invisible to the text distribution. For SigLIP, even Mode 0 is shared with text ($\tilde{\alpha}=0.83$). As representations pass through the LLM, text-aligned variance is amplified (${\sim}150\times$) while MS variance remains flat: MS modes drop from 63.6% of adapter variance to <<1% at the final layer in Ultravox, not destroyed but drowned out.
6 Discussion
6.1 Implications
The fix is objective-side, not encoder-side. If modality collapse is a consequence of the scoring rule being text-shaped, the natural intervention is to reshape the scoring rule. Low-rank adaptation (LoRA; Hu et al., 2022) on the LLM adjusts $q_{\psi}$ to better match $P_{\!M}$, effectively reducing the functional mismatch. Preliminary evidence from concurrent work (Billa, 2026) is consistent: applying LoRA to the LLM backbone reduces text dominance by 23.9%, while adapter-only training increases text dominance by 26.5%. This is predicted by Theorem 2: LoRA reduces the functional mismatch between $P_{\!M}$ and $P_{\!T}$ by reshaping the scoring rule, whereas adapter-only training leaves the decoder’s text-shaped sensitivity unchanged.
However, LoRA on the decoder is necessary but not sufficient: the training objective must explicitly target non-text information. We demonstrate this with two contrasting experiments on Ultravox, using the same LoRA configuration (r=16r{=}16, $\alpha{=}32$, targeting q/k/v/o_proj).
Negative example (text-centric objective). We apply a LoRA checkpoint trained for audio-text conflict resolution (ALME; Billa, 2026) and measure linear probe accuracy on LibriSpeech. Speaker identity accuracy is unchanged (55.0% vs. 54.5% base, within noise), as is lexical accuracy (53.5% vs. 52.6%). The ALME objective rewards resolving text-audio conflicts, not preserving speaker or acoustic attributes, so its gradient signal never reaches modality-specific directions.
Positive example (emotion objective). We train a LoRA on the LLM backbone with a forced-choice emotion detection objective on CREMA-D (6 emotions, 7,442 clips, standard causal LM loss on assistant tokens only). Generation accuracy on held-out samples improves from 17.3% to 61.8% (Table 3). Critically, the emotion probe at llm-final improves from .557.557 to .632.632 (++7.5%), while the speaker probe remains at chance (.135.135 vs. .137.137) and lexical accuracy remains near-ceiling (.994.994). Upstream layers (encoder, adapter) are identical across conditions, confirming the change is entirely in the LLM. The LoRA reshapes the scoring rule so that, at inference, the decoder responds to emotion-relevant directions while leaving others untouched, as Theorem 2 predicts: the emotion objective reduces $W_{\!1}$ along emotion-correlated dimensions without affecting orthogonal ones.
Table 3: LoRA intervention on Ultravox (CREMA-D, llm-final probe accuracy). The emotion LoRA selectively improves emotion accessibility without meaningfully affecting speaker or lexical probes. Generation accuracy measures forced-choice emotion detection on 1,002 held-out samples. The encoder and adapter outputs are identical across conditions (LoRA modifies only the LLM).
Condition Emotion Speaker Lexical Gen. acc.
Base Ultravox .557 .137 1.000 17.3%
+ Emotion LoRA .632 .135 .994 61.8%
$\Delta$ ++7.5% -0.2% -0.6% ++44.5%
Together, the negative and positive examples confirm that the training objective determines which dimensions of the scoring rule are reshaped. A text-centric objective (ALME) leaves the decoder blind to non-text directions; an objective targeting a specific non-text attribute (emotion) teaches the decoder to respond to exactly that dimension.
Text-aligned encoders are a workaround, not a solution. Contrastive encoders like CLIP and SigLIP achieve low mismatch by mapping inputs onto text-correlated dimensions, but this is a projection, not a preservation: modality-specific information with no textual correlate is discarded at the encoder. The Prismatic comparison confirms this: SigLIP’s visual information “improves” through the LLM only because it was pre-projected onto text-like directions. CLIP gives the LLM access to textual correlates of visual information, not visual information itself.
Models must be explicitly trained to use non-text modalities. A text-trained decoder responds only to directions it was trained on (§5.6). No adapter can make it sensitive to directions it has never seen. To build models that genuinely leverage modality-specific information, not just its textual shadow, the training objective must reward extracting non-text directions, whether through LoRA, full multimodal pre-training, or other objective-side interventions. Without this, non-textual information remains hidden in plain sight.
The framework is architecture-agnostic. Our experiments use models with explicit adapter modules, but the analysis does not depend on them. The bound is a property of the decoder’s scoring rule, not of any particular architectural component. Any system where a text-trained decoder processes non-text representations faces the same constraint, whether the representations arrive through a linear projection, an MLP, a Q-Former, a discrete codebook (Borsos et al., 2023; Défossez et al., 2024), or no explicit adapter at all. The LLM’s own layers implicitly serve as the adapter. Discrete-token approaches add a quantization bottleneck that can only increase $W_{\!1}(P_{\!M},P_{\!T})$, making our bound conservative. The core mechanism is the mismatch between the decoder’s training distribution and its inputs at inference, and that mechanism is universal.
6.2 Limitations
Probes as information measures. Linear probes are necessary but not sufficient (Belinkov, 2022), but the key asymmetry (more information is linearly recoverable than the decoder uses) holds regardless. We estimate $L_{\log}$ via p95 gradient norms, not the true supremum. Additional limitations (scope, Assumption 1) are in Appendix D.
[…]
7 Conclusion
[…]
Building multimodal models that truly use what they perceive requires reshaping the decoder’s scoring rule. Text-aligned encoders appear to help but work by discarding non-textual information at the encoder. Our LoRA experiment shows the real fix: training with an emotion objective teaches the decoder to respond to emotion-relevant directions (++7.5% probe, ++44.5% generation accuracy) while leaving others untouched.
The framework extends beyond the models and modalities studied here. The bound is a property of the decoder’s scoring rule, not of any architectural component; no explicit adapter is required. At inference, every trained model has a fixed scoring rule. If that scoring rule was shaped predominantly by text, the constraint applies regardless of architecture or training recipe. The question is not whether modality-specific information can be encoded (it already is) but whether the training objective teaches the decoder to use it. Until it does, that information remains hidden in plain sight.
Reproducibility statement
All models are publicly available. Complete reproduction code, pipeline scripts, and expected-value regression tests are available at https://github.com/jb1999/modality_collapse_paper. Hook points, hyperparameters, and seeds are in Appendix B. All experiments run on a single NVIDIA RTX 3090 Ti (24GB).
Appendix A Theoretical framework: detailed treatment
This appendix provides the full formal development summarized in §4. We aim to make this self-contained: each result is preceded by an intuitive explanation, followed by the formal statement and a detailed proof with annotations.
A.1 Generalized Mutual Information (GMI)
The core question. A representation $Z$ may contain information about a target $Y$ (e.g., the next token). But how much of that information can a specific decoder actually use? Standard mutual information $I(Z;Y)$ answers this for an ideal decoder that is perfectly tuned to the data it receives. A text-trained LLM is not ideal for non-text inputs: it was tuned on text representations, not speech or image representations. We need a measure that accounts for this mismatch. That measure is the Generalized Mutual Information (GMI) (Merhav et al., 1994; Scarlett et al., 2020).
We use the same notation as §3: $q_{\psi}(y|z,c)$ is the decoder’s next-token distribution (Eq. 1), $Z$ is the representation consumed by the LLM, $C$ is the text context, and $Y$ is the target token. Let $\ell_{\psi}(c,z,y):=\log q_{\psi}(y|z,c)$ be the decoder’s log-score (the log-probability it assigns to token $y$ given representation $z$ and context $c$), clipped at a floor $\eta=1/|\mathcal{V}|$ (where $\mathcal{V}$ is the decoder’s output vocabulary) to ensure boundedness.
Intuition. The GMI measures how well the decoder can tell the “right” representation from random alternatives. In words:
[…]
When the decoder reliably scores the correct representation higher than random draws, the GMI is large. When it treats all representations interchangeably (i.e., it cannot tell which $Z$ goes with which $Y$), the GMI collapses to zero. This is closely related to the Donsker-Varadhan variational representation of KL divergence (Donsker & Varadhan, 1983), applied to the decoder’s scoring function rather than an arbitrary test function.
Formally, for a joint distribution $P$ over $(C,Z,Y)$ (either the text distribution $P_{\!T}$ or the modal distribution $P_{\!M}$ from §3), the GMI is:
$\mathrm{GMI}_{P}(q_{\psi})=\mathbb{E}_{(C,Z,Y)\sim P}\!\left[\ell_{\psi}(C,Z,Y)-\log\mathbb{E}_{Z^{\prime}\sim P(\cdot|C,Y)}\!\left[e^{\ell_{\psi}(C,Z^{\prime},Y)}\right]\right].$ (3)
Theorem 1 (Accessible rate = GMI; following Scarlett et al., 2020).
Under i.i.d. sampling from $P_{\!M}$ and standard measurability conditions, the maximum rate extractable by the fixed decoder $q_{\psi}$ is $R_{\mathrm{acc}}(P_{\!M},q_{\psi})=\mathrm{GMI}_{P_{\!M}}(q_{\psi})$.
This is a direct application of the random-coding argument of Scarlett et al. (2020), Chapters 2–3, with constant-composition codebooks under the GMI decoder metric. The key insight is that the decoder $q_{\psi}$ defines a fixed scoring rule (fixed at inference, regardless of how it was trained), and the maximum rate extractable under this rule, averaged over random codebooks drawn from $P_{\!M}$, converges to $\mathrm{GMI}_{P_{\!M}}(q_{\psi})$.
The implication is sharp: regardless of how much information a probe can extract from $Z$, the decoder can never exceed $\mathrm{GMI}_{P_{\!M}}(q_{\psi})$.
A.2 GMI-Wasserstein bound
We now bound the GMI degradation when the decoder sees modal representations $P_{\!M}$ instead of the text representations $P_{\!T}$ it was trained on. The intuition is simple: if modal representations are close to text representations, the decoder should perform nearly as well on them. Two properties of the decoder determine how “close” is close enough: how sensitive it is to small shifts in its input, and how large the space of possible representations is.
Assumption 2 (Local Lipschitz log-score).
There exists $L_{\log}>0$ such that for all $(c,y)$ and all $z,z^{\prime}$ in the typical-state region $\mathcal{Z}$: $|\ell_{\psi}(c,z,y)-\ell_{\psi}(c,z^{\prime},y)|\leq L_{\log}\|z-z^{\prime}\|$.
That is, if you nudge a representation by a small amount, the decoder’s log-probability changes by at most $L_{\log}$ times that amount. The constant $L_{\log}$ captures how sensitive the decoder is to shifts in its input. Neural networks with smooth activations satisfy this locally; we estimate $L_{\log}$ empirically via gradient norms (§5.5).
Assumption 3 (Bounded typical-state region).
The typical-state region $\mathcal{Z}$ has diameter $D=\sup_{z,z^{\prime}\in\mathcal{Z}}\|z-z^{\prime}\|$, which is finite and $O(\sqrt{d})$ after LayerNorm.
This says the representations that actually occur in practice live in a bounded region, not scattered arbitrarily through $\mathbb{R}^{d}$. LayerNorm ensures this by projecting all hidden states onto a hypersphere of radius $\sqrt{d}$.
Theorem 2 (GMI-Wasserstein bound).
Under Assumptions 1, 2, and 3:
$|\mathrm{GMI}_{P_{\!M}}(q_{\psi})-\mathrm{GMI}_{P_{\!T}}(q_{\psi})|\leq\big(1+e^{L_{\log}\cdot D}\big)\cdot L_{\log}\cdot\mathbb{E}_{(C,Y)}\!\left[W_{\!1}\!\big(P_{\!M}(Z|C,Y),\,P_{\!T}(Z|C,Y)\big)\right].$ (4)
Reading the bound. In words:
[…]
The gap between what the decoder could extract from text representations and what it can extract from modal representations is controlled by how far apart those distributions are, amplified by how sensitive the decoder is. Crucially, $W_{\!1}$ measures distributional distance, not just the distance between means: aligning the average modal representation with the average text representation is not enough. If the modal distribution has different shape, spread, or tails, the exponential penalty factor $(1+e^{L_{\log}D})$ catches them. This is why text-aligned encoders reduce the bound (they project onto text-like directions, reducing both the distance and the shape mismatch) while simple mean-centering would not.
The prefactor $(1+e^{L_{\log}D})$ arises from the “competition term” in the proof (see below). In the ambient space this prefactor is vacuously large ($L_{\log}\cdot D$ ranges from 10 to 150); the support-restricted bound (§A.3) tightens it using the effective support diameter.
A.2.1 Proof of Theorem 2
Proof overview. The GMI has two pieces: (A) how well the decoder scores the correct representation, and (B) how well it scores random alternatives. We need to show that switching from text to modal representations does not change either piece by too much. For piece (A), the argument is simple: if the decoder’s scores change smoothly (Lipschitz assumption), then moving representations a small distance changes expected scores by a small amount. For piece (B), the argument is harder because piece (B) involves an exponential of the scores, which amplifies small changes. The bounded-diameter assumption (representations live in a finite region) keeps this amplification under control.
Setup. Write $\ell(z)\equiv\ell_{\psi}(c,z,y)$ for fixed $(c,y)$. The GMI functional decomposes as $\Gamma(P)=A(P)-B(P)$ where:
$A(P)$ $=\mathbb{E}_{Z\sim P}[\ell(Z)]$ (direct term: expected score of the correct representation) (5)
$B(P)$ $=\log\mathbb{E}_{Z\sim P}[e^{\ell(Z)}]$ (competition term: how well random alternatives score) (6)
We bound $|A(P_{\!M})-A(P_{\!T})|$ and $|B(P_{\!M})-B(P_{\!T})|$ separately.
Step 1: Direct term (how much does the correct representation’s score change?). We want to bound how much the expected score changes when we switch from $P_{\!T}$ to $P_{\!M}$. The Kantorovich–Rubinstein theorem (Kantorovich & Rubinstein, 1958; Villani, 2009) says: if a function changes by at most $L$ when you move its input by 1 unit ($L$-Lipschitz), then the difference in its expected value under two distributions is at most $L$ times the Wasserstein distance between those distributions. Since $\ell$ is $L_{\log}$-Lipschitz (Assumption 2):
$|A(P_{\!M})-A(P_{\!T})|=|\mathbb{E}_{P_{\!M}}[\ell]-\mathbb{E}_{P_{\!T}}[\ell]|\leq L_{\log}\cdot W_{\!1}(P_{\!M},P_{\!T}).$ (7)
In words: if modal and text representations are close (small $W_{\!1}$) and the decoder’s scores change smoothly (small $L_{\log}$), then the expected score barely changes.
Step 2: Competition term (how much does the random-alternative score change?). This is harder because $B(P)=\log\mathbb{E}[e^{\ell(Z)}]$ involves an exponential of the score, which amplifies small differences. A coupling $\pi$ is a way of pairing up samples from $P_{\!M}$ and $P_{\!T}$ (imagine lining up modal representations next to the text representations they most resemble). For each pair $(Z_{M},Z_{T})$, define $\Delta\ell=\ell(Z_{M})-\ell(Z_{T})$: the difference in score between the paired representations. By the Lipschitz assumption, $|\Delta\ell|\leq L_{\log}\cdot\|Z_{M}-Z_{T}\|$, and by the bounded diameter (Assumption 3), $|\Delta\ell|\leq L_{\log}\cdot D$.
Write $m(P)=\mathbb{E}_{Z\sim P}[e^{\ell(Z)}]$. Then:
$\frac{m(P_{\!M})}{m(P_{\!T})}$ $=\mathbb{E}_{\pi}\!\left[\frac{e^{\ell(Z_{M})}}{m(P_{\!T})}\right]=\mathbb{E}_{\pi}\!\left[\frac{e^{\ell(Z_{T})}}{m(P_{\!T})}\cdot e^{\Delta\ell}\right].$ (8)
The first factor $e^{\ell(Z_{T})}/m(P_{\!T})$ is a probability weight under $P_{\!T}$ (it sums to 1). The second factor $e^{\Delta\ell}$ captures how the shift from $Z_{T}$ to $Z_{M}$ changes the score. Since $|e^{x}-1|\leq|x|\cdot e^{|x|}$ for bounded $|x|\leq L_{\log}D$:
$\left|\frac{m(P_{\!M})}{m(P_{\!T})}-1\right|\leq e^{L_{\log}D}\cdot\mathbb{E}_{\pi}[|\Delta\ell|]\leq e^{L_{\log}D}\cdot L_{\log}\cdot\mathbb{E}_{\pi}[\|Z_{M}-Z_{T}\|].$ (9)
Using $|\log(1+x)|\leq|x|$ for $|x|<1$ (applicable when $W_{\!1}$ is small enough that the ratio stays near 1):
$|B(P_{\!M})-B(P_{\!T})|=\left|\log\frac{m(P_{\!M})}{m(P_{\!T})}\right|\leq e^{L_{\log}D}\cdot L_{\log}\cdot W_{\!1}.$ (10)
The $e^{L_{\log}D}$ prefactor is the price of the exponential: the competition term is more sensitive to distributional shift than the direct term because it involves $e^{\ell}$ rather than $\ell$. This is analogous to how compound interest amplifies small rate changes: a 1% shift in the base rate produces a much larger change in the compounded amount when the compounding period is long.
Step 3: Combine. The total GMI change is bounded by the sum of the direct and competition terms:
$|\Gamma(P_{\!M})-\Gamma(P_{\!T})|\leq\underbrace{L_{\log}\cdot W_{\!1}}_{\text{direct}}+\underbrace{e^{L_{\log}D}\cdot L_{\log}\cdot W_{\!1}}_{\text{competition}}=(1+e^{L_{\log}D})\cdot L_{\log}\cdot W_{\!1}.$ (11)
The competition term dominates: it is $e^{L_{\log}D}$ times larger than the direct term. This is why the bound’s prefactor is large, and why the support-restricted corollary (which reduces $D$) is practically important.
Step 4: Average over $(C,Y)$. Everything above was for a fixed prompt $C$ and target $Y$. By Assumption 1 ($P_{\!M}(C,Y)=P_{\!T}(C,Y)$), the same prompts and targets appear under both laws, so we simply average over all $(C,Y)$ pairs to obtain the final bound.
A.3 Support-restricted bound
Support-restricted bound. If $P_{\!M}(Z|c,y)$ and $P_{\!T}(Z|c,y)$ are both supported within a region $\mathcal{S}\subseteq\mathcal{Z}$ with $D_{\mathcal{S}}=\mathrm{diam}(\mathcal{S})\leq D$, then $D$ may be replaced by $D_{\mathcal{S}}$ in Theorem 2.
Proof. If both marginals are supported in $\mathcal{S}$, any coupling $\pi$ has support on $\mathcal{S}\times\mathcal{S}$, so $\|Z_{M}-Z_{T}\|\leq D_{\mathcal{S}}$ for $\pi$-a.e. $(Z_{M},Z_{T})$. Step 2 above then bounds $|\Delta\ell|\leq L_{\log}\cdot D_{\mathcal{S}}$, giving the tighter prefactor.
Why this matters. The ambient diameter $D$ makes the prefactor $(1+e^{L_{\log}D})$ astronomically large ($L_{\log}\cdot D$ ranges from 10 to 150). But modal and text representations often occupy a much smaller effective region. We estimate the effective support diameter $D_{\text{eff}}$ via the participation ratio of the pooled representation covariance (§5.5), yielding $L_{\log}\cdot D_{\text{eff}}\in[3,7]$ and a moderate prefactor ($\sim$20–1100).
A.4 Probe-decoder asymmetry
Theorem 3 (Probe penalty).
Let $h(a|z)$ be a linear probe with $\log h$ being $L_{h}$-Lipschitz in $z$. Then: $|\mathbb{E}_{P_{\!M}}[\log h(A|Z)]-\mathbb{E}_{P_{\!T}}[\log h(A|Z)]|\leq L_{h}\cdot W_{\!1}(P_{\!M},P_{\!T})$.
Proof. Identical to Step 1 of the Theorem 2 proof, replacing $\ell_{\psi}$ with $\log h$ and $L_{\log}$ with $L_{h}$. Crucially, there is no competition term for probes (no random-coding argument), so the exponential prefactor does not appear.
The key asymmetry. A linear probe is a simple instrument: one matrix multiplication followed by a softmax. Its sensitivity to distributional shift ($L_{h}$) is just the size of that matrix’s weights. The LLM decoder is a deep transformer with billions of parameters, and its sensitivity ($L_{\log}$) reflects that depth. Empirically, $L_{\log}/L_{h}\approx 30\times$5.5).
Now consider the same distributional shift (modal representations replacing text representations). The probe’s bound says degradation $\leq L_{h}\cdot W_{\!1}$ (small). The decoder’s bound says degradation $\leq(1+e^{L_{\log}D})\cdot L_{\log}\cdot W_{\!1}$ (large). Same shift, dramatically different impact. This is the formal basis for the “present but inaccessible” phenomenon: a probe can still read the information because it is a simple, low-sensitivity instrument, but the decoder’s complex scoring rule is far more sensitive to the distributional mismatch and cannot use it.
Appendix B Experimental details
B.1 Hook points and probing hyperparameters
Model Encoder Adapter LLM-16 LLM-final
Ultravox audio_tower.layer_norm multi_modal_projector.ln_post language_model.model.layers.16 language_model.model.norm
Qwen2-Audio audio_tower.layer_norm multi_modal_projector language_model.model.layers.16 language_model.model.norm
LLaVA ...vision_tower. ...post_layernorm model.mm_projector ...language_model. layers.16 ...language_model. norm
Table 4: Hook points for representation extraction.
Logistic regression with $\ell_{2}$ regularization ($C=1.0$, scikit-learn default). Five random seeds (42, 43, 44, 45, 46). 80/20 stratified train/test split. Input features $z$-score normalized. Mean $\pm$ std reported.
B.2 Lipschitz estimation
$L_{\log}$ estimated via the 2-nearest-neighbor gradient norm method over n=1,000n=1{,}000 samples. For each sample $z_{i}$, we compute $\|\nabla_{z}\log q_{\psi}(y_{i}|z_{i},c_{i})\|$ via autograd. The 95th percentile of the resulting distribution serves as the $L_{\log}$ estimate.
B.3 Extraction prompts
During representation extraction, each model receives a minimal, task-neutral prompt to avoid biasing hidden states toward any particular downstream task. Representations are extracted from intermediate hook points, not from the generated output.
Speech models. Ultravox receives the audio placeholder token <|audio|> wrapped in the model’s chat template (system/user/assistant turns). Qwen2-Audio requires a text prompt alongside audio; we use "Describe this audio." in the user turn.
[…]
Text baselines. Text baselines (Llama-3.1-8B for speech models, Vicuna-7B for vision models) process raw transcripts or captions with no additional prompt, using the model’s default tokenization.
B.4 LoRA experiment details
The emotion LoRA (§6.1) uses standard causal LM fine-tuning on Ultravox with a forced-choice emotion detection objective on CREMA-D.
LoRA configuration. Rank r=16r{=}16, $\alpha{=}32$, dropout $0.05$, target modules: q_proj, k_proj, v_proj, o_proj (same configuration as the ALME LoRA used as the negative example).
Training. AdamW optimizer, learning rate $10^{-4}$, weight decay $0.01$, gradient accumulation $8$ steps (effective batch size $8$), gradient clipping at max norm $1.0$, mixed precision (bfloat16), gradient checkpointing enabled. Five epochs with early stopping (patience $2$, monitoring validation loss). Best checkpoint at epoch 5 (val loss $0.0842$, token accuracy $96.7\%$).
Prompt template. Each training sample uses a three-turn chat format:
  • System: “You are an expert speech analyst. When presented with audio, identify the emotion expressed by the speaker. Always prioritize what you HEAR in the audio.”
[…]
  • Assistant: {"answer": "<emotion>"} (training target)
Loss is computed only on the assistant response tokens; all input/prompt tokens are masked with label 100-100.
Data split. CREMA-D (7,442 clips, 91 speakers, 6 emotions). 80/20 stratified split yields 1,002 held-out evaluation samples (167 per emotion class), saved deterministically for reproducibility.
Baseline evaluation. Base Ultravox (no LoRA) achieves $17.3\%$ generation accuracy on the held-out split (near chance for 6 classes), defaulting to “neutral” for most inputs.
B.5 Full probe accuracy tables
Table 5 reports the full probe accuracies (mean $\pm$ std over 5 seeds) at all four hook points for every model–dataset–information-type combination. The key pattern to observe: probe accuracy for non-text information types remains well above chance even at the LLM’s final layer, confirming that the information survives the decoder’s processing. It is the decoder’s use of this information, not its presence, that fails.
Model Info type (Dataset) Encoder Adapter LLM-16 LLM-Final
Ultravox Lexical (LS) $422\pm.014 .$$279\pm.018 .$$570\pm.023 .$$535\pm.027$
Speaker (LS) $721\pm.029 .$$605\pm.031 .$$628\pm.019 .$$556\pm.019$
Emotion (CD) $679\pm.012 .$$621\pm.007 .$$603\pm.011 .$$557\pm.006$
Speaker (CD) $508\pm.005 .$$196\pm.010 .$$204\pm.011 .$$137\pm.008$
Acoustic (E50) $760\pm.033 .$$647\pm.041 .$$645\pm.035 .$$581\pm.032$
Q2-Audio Lexical (LS) $270\pm.006 .$$242\pm.015 .$$506\pm.018 .$$471\pm.021$
Speaker (LS) $980\pm.004 .$$961\pm.006 .$$665\pm.022 .$$589\pm.022$
Emotion (CD) $866\pm.007 .$$74\pm.010 .88$$80\pm.006$ $877\pm.004$
Speaker (CD) $880\pm.009 .$$816\pm.003 .$$725\pm.016 .$$556\pm.012$
Acoustic (E50) $955\pm.005 .$$946\pm.011 .$$962\pm.004 .$$965\pm.003$
LLaVA Lexical (CO) $629\pm.011 .$$620\pm.010 .$$75\pm.004 .66$$69\pm.006$
Object cat. (CO) $84\pm.011 .77$$70\pm.009$ $793\pm.010 .$$808\pm.011$
Super cat. (CO) $802\pm.011 .$$797\pm.007 .$$833\pm.009 .$$842\pm.012$
Prism.-D Lexical (CO) $608\pm.007 .$$637\pm.011 .$$52\pm.011 .66$$64\pm.011$
Object cat. (CO) $795\pm.016 .$$792\pm.015 .$$794\pm.014 .$$792\pm.012$
Super cat. (CO) $837\pm.013 .$$832\pm.010 .$$834\pm.011 .$$839\pm.014$
Obj. count (CO) $581\pm.017 .$$594\pm.015 .$$611\pm.015 .$$614\pm.016$
Obj. size (CO) $667\pm.006 .$$679\pm.010 .$$693\pm.010 .$$688\pm.007$
Spatial spr. (CO) $470\pm.013 .$$465\pm.025 .$$475\pm.010 .$$487\pm.008$
Prism.-S Lexical (CO) $638\pm.007 .$$657\pm.008 .$$79\pm.007 .66$$63\pm.009$
Object cat. (CO) $813\pm.007 .$$804\pm.011 .$$811\pm.010 .$$819\pm.014$
Super cat. (CO) $837\pm.013 .$$832\pm.017 .$$849\pm.010 .$$851\pm.010$
Obj. count (CO) $594\pm.003 .$$598\pm.013 .$$652\pm.007 .$$647\pm.014$
Obj. size (CO) $54\pm.019 .66$$69\pm.014$ $722\pm.018 .$$699\pm.020$
Spatial spr. (CO) $478\pm.021 .$$459\pm.015 .$$507\pm.014 .$$481\pm.023$
Table 5: Full probe accuracy (mean $\pm$ std over 5 seeds) for all models, datasets, hook points, and information types.
B.6 Information retention overview
Table 6 summarizes the information retention pattern across all five models. Retention is computed as the ratio of probe accuracy at the LLM’s final layer to probe accuracy at the adapter output, normalized by chance. Values above 100% indicate the LLM amplifies the signal; values below 100% indicate degradation. The key pattern: non-aligned encoders show asymmetric behavior (lexical content is amplified while speaker identity collapses), whereas text-aligned encoders retain or slightly improve all information types, but only because the encoder has already discarded non-textual information.
Table 6: Information retention (adapter $\to$ LLM-final, %). Values >100{>}100% indicate the LLM amplifies information; <100{<}100% indicates degradation. Text-aligned encoders retain all types; non-aligned encoders amplify lexical content while speaker identity collapses by up to 39%. LS = LibriSpeech, CD = CREMA-D, CO = COCO.
Model Encoder Lexical Best non-text Worst non-text
Encoder not text-aligned
Ultravox Whisper 192% 92% (speaker, LS) 70% (speaker, CD)
Qwen2-Audio Whisper 195% 102% (acoustic) 61% (speaker, LS)
Prismatic-D DINOv2 104% 101% (super cat.) 100% (object cat.)
Encoder text-aligned
LLaVA CLIP 108% 106% (super cat.) 105% (object cat.)
Prismatic-S SigLIP 101% 102% (obj. & super cat.) 102% (obj. & super cat.)
B.7 Detailed information retention
Table 7 provides the full breakdown of information retention per model and information type, with probe accuracies at the adapter and LLM-final layers.
Model Info type (Dataset) Adapter LLM-Final Retention (%)
Encoder not text-aligned
Ultravox Lexical (LS) .279 .535 192
Speaker (LS) .605 .556 92
Emotion (CD) .621 .557 90
Speaker (CD) .196 .137 70
Acoustic (E50) .647 .581 90
Q2-Audio Lexical (LS) .242 .471 195
Speaker (LS) .961 .589 61
Emotion (CD) .874 .877 100
Speaker (CD) .816 .556 68
Acoustic (E50) .946 .965 102
Prism.-D Lexical (CO) .637 .664 104
Object cat. (CO) .792 .792 100
Super cat. (CO) .832 .839 101
Encoder text-aligned
LLaVA Lexical (CO) .620 .669 108
Object cat. (CO) .770 .808 105
Super cat. (CO) .797 .842 106
Prism.-S Lexical (CO) .657 .663 101
Object cat. (CO) .804 .819 102
Super cat. (CO) .832 .851 102
Table 7: Information retention from adapter to LLM-final layer (probe accuracy ratio $\times$ 100). Retention >{>}100% means the LLM amplifies the signal; <{<}100% means degradation. Non-aligned encoders show selective amplification (lexical recovers, speaker degrades); text-aligned encoders retain or slightly improve all types. LS = LibriSpeech, CD = CREMA-D, E50 = ESC-50, CO = COCO.
B.8 Lipschitz constants and Wasserstein distances
Table 8 reports the empirical Lipschitz constants and Wasserstein distances used to evaluate the bound in §5.4.
Model Dataset $L_{\log}$ $L_{\log}$ (p95) $D$ $W_{\!1}$ $L_{\log}\cdot W_{\!1}$
Speech
Ultravox LibriSpeech 9.08 14.58 10.3 17.8 162
CREMA-D 5.82 9.70 17.0
ESC-50 3.66 5.92 20.7
Q2-Audio LibriSpeech 0.49 0.67 122.2 38.9 19.1
CREMA-D 0.59 0.75 229.9
ESC-50 0.47 0.59 202.5
Vision
LLaVA COCO 0.29 0.36 28.5 46.3 13.4
Prism.-D COCO 0.40 0.49 35.7 50.4 20.2
Prism.-S COCO 1.12 1.44 16.2 47.9 53.7
Table 8: Lipschitz constants ($L_{\log}$, mean and p95 of gradient norm distribution), representation diameter ($D$), Wasserstein-1 distance ($W_{\!1}$ at LLM layer 16), and their product. $L_{\log}$ is content-dependent: it decreases as audio moves away from text (LibriSpeech >> CREMA-D >> ESC-50). $L_{\log}\cdot W_{\!1}$ is largest for Ultravox (162, strongest degradation) and smallest for LLaVA (13.4, no degradation). Prismatic-S has $2.8\times$ higher $L_{\log}$ than Prismatic-D, consistent with the LLM being more responsive to text-aligned representations.
B.9 Audio-token-only pooling
Our standard extraction mean-pools LLM-layer activations over all sequence positions (system prompt, audio tokens, assistant prompt). This could dilute the audio signal. Ultravox’s processor exposes the audio token boundaries (audio_token_start_idx, audio_token_len), letting us restrict pooling to audio positions only. We re-extracted LLM representations with audio-only pooling and re-ran probes to compare.
Table 9 shows the result: audio-only pooling produces nearly identical probe accuracies across all datasets and information types, with all deltas within $\pm$2.1% and most under 1%. Restricting to audio tokens does not help, and in several cases slightly hurts. This is because the LLM’s self-attention layers redistribute information across all sequence positions as they process the input. By layer 16, audio-derived information has spread to text prompt tokens as well. Pooling over the full sequence therefore captures more of the available signal, not less. This validates our use of full-sequence pooling throughout the paper: it is not a conservative approximation that dilutes the audio signal, but rather the appropriate strategy given how transformers distribute information.
Dataset Label All-token Audio-only $\Delta$
LLM hidden layer 16
[…]
LLM final layer
[…]
Table 9: Probe accuracy with all-token vs. audio-only pooling at LLM layers (Ultravox). All deltas are within $\pm$2.1%, most under 1%. Audio-only pooling does not improve probe accuracy. The LLM’s self-attention layers mix information across all sequence positions: by layer 16, audio-derived information has been redistributed to text prompt tokens as well. Restricting pooling to audio positions therefore discards information that the LLM has already spread across the full sequence, slightly reducing the available signal. This validates our use of full-sequence pooling throughout the paper.
Appendix C Additional figures
Figure 1: Probe accuracy trajectories for all five models. Rows 1–2: speech models (Ultravox, Qwen2-Audio) across LibriSpeech, CREMA-D, and ESC-50. Row 3: vision models (LLaVA, Prismatic-DINOv2, Prismatic-SigLIP) on COCO.
Figure 2: Mode alignment profiles for all five models. Blue: alignment score $\tilde{\alpha}(u_{k})$; orange: eigenvalue spectrum $\lambda_{k}$ (log scale). The dominant eigenmodes are modality-specific ($\tilde{\alpha}\approx 0$) for all models with non-text-aligned encoders; for Prismatic-S (SigLIP, text-aligned), even Mode 0 is text-aligned ($\tilde{\alpha}=0.83$).
The Prismatic comparison (Table 10) isolates encoder text-alignment as the causal variable. Both VLMs share the same architecture, adapter, training recipe, and LLM backbone; only the vision encoder differs. SigLIP (text-aligned) shows consistent improvement through the LLM for all information types, while DINOv2 (not text-aligned) shows stagnation or minimal change for non-textual attributes.
Probe accuracy
Encoder Info type Adapter LLM-final $\Delta_{\text{LLM}}$
Text-describable attributes
DINOv2 Lexical .637 .664 +4.2%
Object cat. .792 .792 -0.0%
Super cat. .832 .839 +0.8%
SigLIP Lexical .657 .663 +0.9%
Object cat. .804 .819 +1.9%
Super cat. .832 .851 +2.3%
Non-textual attributes
DINOv2 Obj. count .594 .614 +3.4%
Obj. size .679 .688 +1.3%
Spatial spr. .465 .487 +4.7%
SigLIP Obj. count .598 .647 +8.2%
Obj. size .669 .699 +4.5%
Spatial spr. .459 .481 +4.8%
Table 10: Prismatic controlled comparison (full results). $\Delta_{\text{LLM}}$ = change from adapter to LLM-final. Top: text-describable attributes (named in captions). Bottom: non-textual attributes (object count, average object size, spatial spread; derived from annotations, not captions). Text baseline accuracy is substantially lower (.415–.495) for non-textual attributes, confirming their visual nature. Same architecture, same LLM, only encoder differs.
The mode alignment spectrum (Table 11) reveals which directions carry the mismatch. For each model, we decompose the adapter output covariance into eigenmodes and measure their text alignment. The dominant eigenmode is modality-specific ($\tilde{\alpha}\ll 1$) for all non-text-aligned encoders, while for SigLIP even the dominant mode is text-aligned ($\tilde{\alpha}=0.83$).
Model Mode 0 $\tilde{\alpha}$ Mean $\tilde{\alpha}$ (top 10) MS modes MS var.
Ultravox 0.001 1.76 11/100 63.6%
Qwen2-Audio 0.010 7.21 5/100 97.6%
LLaVA 0.013 3.94 9/100 71.9%
Prismatic-D 0.034 0.087 53/100 71.0%
Prismatic-S 0.827 1.95 14/100 24.6%
Table 11: Mode alignment summary. MS modes = modes with $\tilde{\alpha}<0.5$ (modality-specific). MS var. = fraction of top-100 eigenmode variance in MS modes. Mode 0 is modality-specific ($\tilde{\alpha}\ll 1$) for all non-text-aligned encoders; for SigLIP (text-aligned), even Mode 0 is text-aligned ($\tilde{\alpha}=0.83$). Extending to all 4096 modes strengthens the pattern: tail modes are overwhelmingly text-aligned with negligible individual variance, so the MS/TA split is driven by the high-variance modes reported here.
Appendix D Additional limitations
Scope. We test speech and vision across five models. The framework applies to any multimodal LLM with a text-trained decoder, but video/3D remain untested.
Assumption 1 (shared marginal). We assume $P_{\!M}(C,Y)=P_{\!T}(C,Y)$. This holds when the multimodal model is trained on text-centric tasks (transcription, captioning). Violations add a marginal-shift term independent of representation geometry.
Generality of the GMI constraint. The mechanism described in the main text is not architectural: any multimodal LLM whose decoder was trained under a text-dominated objective has a text-shaped scoring rule, regardless of whether the adapter is a linear projection, MLP, Q-Former, perceiver resampler, or discrete codebook. The GMI constraint applies to multimodal modelling in general: unless the training objective explicitly targets each modality, the decoder remains indifferent to non-text directions.

### v2

@@
 Abstract
-[…]
-We formalize this as a mismatched decoder problem: a decoder trained on text can only extract information along text-aligned directions.
-Accessible information is bounded by the Generalized Mutual Information (GMI), with degradation scaling with distributional distance and decoder sensitivity.
-The bound is a property of the decoder’s scoring rule, not of any particular architecture; it applies whether non-text inputs arrive through a learned projection, a discrete codebook, or no explicit adapter at all.
-We validate this across five models spanning speech and vision.
-A controlled experiment (two Prismatic VLMs differing only in encoder text-alignment) confirms the bottleneck is the decoder’s scoring rule, not the encoder or projection.
-A LoRA intervention demonstrates the fix: training with an emotion objective improves emotion accessibility (++7.5%) without affecting other attributes, confirming that the training objective determines what becomes accessible.
+Numerous studies have shown that multimodal LLMs process speech and images well but fail in non-intuitive ways rendering trivial tasks such as object counting unreliable.
+We investigate this behavior from an information-theoretic perspective by framing multimodal LLM inference as a mismatched decoder problem: a decoder trained primarily on text can only extract information along text-aligned directions (removing up to 98% of the variation in modality-specific (non-text) directions improves decoder loss) and the amount of accessible information is bounded by the Generalized Mutual Information (GMI).
+We show that information loss is bounded as the distributional mismatch between the source data and the text data increases, and as the sensitivity of the decoder increases.
+This bound is a function of the model’s scoring rule not its architecture.
+We validate the predictions across five models spanning speech and vision.
+A controlled study (two Prismatic VLMs differing only in encoder text-alignment) shows that the bottleneck lies in the scoring rule of the decoder rather than the text-alignment of the encoder or the learned projection.
+A LoRA intervention demonstrates that simply training with an emotion-related objective improves emotion detection from 17.3% to 61.8% task accuracy without affecting other attributes, confirming that the training objective determines what becomes accessible.
 1 Introduction
-Multimodal LLMs follow a common recipe: an encoder processes the non-text input (speech, image), a learned projection maps it into the LLM’s embedding space, and the LLM generates text (Liu et al., 2023; Dai et al., 2023; Fixie AI, 2024; Chu et al., 2024; Karamcheti et al., 2024).
-By standard benchmarks these systems appear to work well: speech models transcribe accurately, vision models caption faithfully.
-But this success is a mirage.
-These systems excel at tasks requiring only text-correlated information (the words spoken, the objects visible) while silently discarding everything else.
-A speech model that transcribes perfectly may have no access to who is speaking or how they feel; a vision model that names every object may be blind to spatial layout or texture.
-This paper shows that the loss is structural and predictable.
-[…]
-Worse, modality-specific information degrades decoding.
-Removing 64% of the variance in modality-specific directions improves cross-entropy loss; the decoder works better with less information.
-Lexical probe accuracy nearly doubles from the adapter to the final LLM layer (the decoder amplifies what it knows how to use), while speaker probe accuracy drops by up to 39% over the same path.
-The mechanism is not active suppression: the decoder simply has no learned response to non-text directions.
-It is indifferent to them, but that indifference has consequences, because the decoder treats unfamiliar variance as noise.
-Where the collapse manifests depends on the encoder: non-aligned encoders (Whisper, DINOv2) pass rich representations the decoder is blind to; text-aligned encoders (CLIP, SigLIP) discard non-textual information before it reaches the decoder.
-Either way, the information is lost (Appendix Table 6).
-Our thesis.
-We formalize this as a mismatched decoder problem (Merhav et al., 1994; Scarlett et al., 2020).
-The LLM’s scoring rule was learned from text, so it can only extract information along text-aligned directions.
-Accessible information is bounded by the Generalized Mutual Information (GMI), not the standard mutual information.
-We call the difference the information accessibility gap.
+Multimodal LLMs follow a common recipe: an encoder processes the non-text input (speech, image), a learned projection maps it into the LLM’s embedding space, and the LLM generates a text response (Liu et al., 2023; Dai et al., 2023; Fixie AI, 2024; Chu et al., 2024; Karamcheti et al., 2024).
+These systems appear to perform well on standard benchmarks.
+However, there are many reported gaps in model performance (Chen et al., 2026; Cuervo et al., 2026; Thrush et al., 2022; Yüksekgönül et al., 2023; Jain et al., 2025) where even some trivial tasks are not resolved.
+A model might produce a perfect response to speech input but have no understanding of the emotional state of the speaker; it may correctly identify every object in an image and yet fail to count the number of objects or understand the spatial relationships.
+Here, the information is clearly available to the model, since it can list every object, but it cannot make use of this available information for all tasks.
+We refer to this selective failure as modality collapse.
+To provide insight into this behavior we borrow the well-defined concept of a mismatched decoder from communication theory, where a decoder originally designed for source A receives source B.
+A multimodal LLM fits this definition of a mismatched decoder exactly: an LLM trained to ”decode” text representation is now expected to extract information from audio and/or image projections.
+Our mismatched decoder analysis is a formalization of the perception constraint cost (Blau and Michaeli, 2019).
+In this framing the accessible information is bounded by the Generalized Mutual Information (GMI) not standard mutual information, as would be the case for matched decoders.
+We refer to this difference as the information accessibility gap.
 The gap is not about architecture; it is about the scoring rule.
@@
 What matters is the training objective, not which components were updated (§6.1).
-What determines the size of the gap?
-A linear probe simply measures whether the information it was trained to detect is present.
-The decoder is a different instrument entirely.
-It is a deep transformer trained on text, and it is highly sensitive to its input, responding to variance in every direction, including non-text directions (we confirm this empirically via gradient isotropy in §5.5).
-But its response to non-text variance is unproductive: the decoder interprets unfamiliar structure as noise, and that noise disrupts its text processing.
-This is why removing modality-specific directions helps: not because the decoder ignores them, but because it responds to them in ways that degrade output.
-We formalize this: degradation scales with both the distributional distance between modal and text representations and the decoder’s sensitivity to that distance (Theorem 2), and the decoder’s sensitivity is up to $30\times$ that of a probe (Theorem 3, §A.4).
-Text-aligned encoders reduce the distance (§5.3), but only by discarding non-textual information upstream.
-The collapse happens earlier, not less.
+What causes this gap?
+A transformer model is sensitive to its input in all directions; one trained on text is still sensitive to non-text directions (we validate this empirically using gradient isotropy in §5.6).
+That said, its sensitivity to non-text directions is unproductive.
+The decoder treats unfamiliar structure as noise, and noise hurts its text processing, as would be expected in a mismatched decoder.
+In §5.4 we empirically show that removing up to 98% of the variation in modality-specific (non-text) directions improves decoder loss; the decoder is not indifferent to non-text directions, but responds to them in destructive ways.
+We can quantify this degradation (Theorem 2).
+It is proportional to both the distributional distance between the modality and text representations and the decoder’s sensitivity to this distance (its Lipschitz constant $L_{\log}$, which measures how much the output changes per unit change in the input).
+This sensitivity can be up to $30\times$ that of a linear probe (a classifier trained on frozen representations to test what information is present; Theorem 3, §A.5).
+Text-aligned encoders reduce the distance (§5.5), but only by discarding non-textual information earlier on.
+The collapse occurs earlier, not less.
 Contributions.
-(1) We formalize modality collapse as mismatched decoding and prove that accessible information is bounded by the GMI, with degradation governed by distributional distance and decoder sensitivity (§4).
-(2) Across five models and two modalities, we demonstrate the information accessibility gap: non-text information survives through the LLM but degrades decoding because the decoder has no learned use for it (§5.2, §5.4).
-(3) A controlled experiment with Prismatic VLMs (same architecture, same LLM, only encoder text-alignment differs) isolates the scoring rule as the causal variable (§5.3).
-(4) A LoRA intervention confirms the prescription: training with an emotion objective teaches the decoder to respond to that dimension (++7.5%), while a text-centric objective leaves it indifferent (§6.1).
+(1) We formalize modality collapse as mismatched decoding, and prove that the accessible information is at most the GMI, and degrades with the distributional mismatch and the sensitivity of the decoder (§4).
+(2) On five models and two modalities, we show that the information accessibility gap is present: non-text information is preserved by the LLM but is not decodable because the decoder is not incentivized to use it (§5.4, §5.2).
+(3) We run a controlled study on Prismatic VLMs (identical architecture, identical LLM, only text-alignment of the encoder changes), and show that the scoring rule is the causal factor (§5.5).
+(4) We run a LoRA intervention that confirms the prescription: training with an emotion objective improves emotion task accuracy from 17.3% to 61.8% (++7.5% probe), while training with a text objective does not (§6.1).
 2 Related work
 Multimodal LLMs and the adapter paradigm.
-The strategy of projecting encoded features into an LLM’s embedding space has become standard for both vision (Liu et al., 2023; Dai et al., 2023; Zhu et al., 2024; Karamcheti et al., 2024; Tong et al., 2024) and speech (Fixie AI, 2024; Chu et al., 2024).
-These systems achieve strong performance on text-centric benchmarks but struggle with tasks requiring modality-specific information, an observation that has been noted empirically but not explained theoretically.
+The practice of projecting encoded representations into the embedding space of an LLM has been widely adopted for both vision (Liu et al., 2023; Dai et al., 2023; Zhu et al., 2024; Karamcheti et al., 2024; Tong et al., 2024) and speech (Fixie AI, 2024; Chu et al., 2024).
+While these approaches perform well on text-centric tasks, they perform poorly on tasks that require modality-specific knowledge, a phenomenon observed empirically but not understood theoretically.
 Modality gap.
 Liang et al.
-(2022) showed that CLIP embeddings of different modalities occupy distinct regions of the shared space, separated by a persistent “modality gap.” Huh et al.
-(2024) observed convergence of representations across modalities.
-Our work complements these geometric observations with an information-theoretic mechanism: the gap matters not because information is destroyed, but because the decoder cannot access information in unfamiliar directions.
+(2022) observed that the CLIP embeddings of different modalities are separated by a “modality gap” in the shared space.
+More recently, Huh et al.
+(2024) reported that the representations of different modalities become similar.
+While our results are largely orthogonal to these geometric findings, we provide an information-theoretic explanation for the modality gap: it is not that information is lost, but rather that the decoder is unable to make use of information in unfamiliar directions.
+Modality collapse.
+Javaloy et al.
+(2022) use this term to describe models that discard an entire modality during training due to gradient conflicts.
+Our usage describes a subtler failure: the model processes all modalities and preserves their information internally, but the decoder selectively fails to use non-primary-modality content.
 Mismatched decoding.
-The theory of mismatched decoders (Merhav et al., 1994; Lapidoth, 1996; Scarlett et al., 2020) characterizes the information rate achievable when a decoder is designed for a different channel than the one actually used.
-We import this framework into the multimodal LLM setting: at inference, any LLM whose scoring rule was shaped by text-dominated training is a mismatched decoder for non-text inputs.
-The mismatch is determined by the training objective, not by which components were updated.
+The mismatched decoder framework (Merhav et al., 1994; Lapidoth, 1996; Scarlett et al., 2020) provides the data rates supported by a decoder mismatched to its incoming channel.
+We borrow this approach and apply it to the multimodal LLM framework.
+At inference time, any LLM that uses a scoring rule that was learned via a text-dominant training procedure is a mismatched decoder when applied to non-text inputs.
+The mismatch arises from the loss function, not from what parameters were updated.
 Representation probing.
-Linear probes are a standard tool for measuring information content in neural representations (Alain & Bengio, 2017; Conneau et al., 2018; Hewitt & Manning, 2019; Belinkov, 2022).
-We use probes not to claim information is “encoded” in a representation, but to demonstrate the gap between information present (probe-recoverable) and information accessible (decoder-extractable).
-Rate-distortion in ML.
-Blau & Michaeli (2019) introduced the rate-distortion-perception tradeoff, showing that perceptual quality constraints can increase the rate needed for lossy compression.
-Our theoretical framework shares the insight that the output distribution shape (here, conformity to the text manifold) imposes additional costs beyond standard rate-distortion.
+Linear probes are a common method for probing the information content of neural representations (Alain and Bengio, 2017; Conneau et al., 2018; Hewitt and Manning, 2019; Belinkov, 2022).
+We do not use probing to argue that some information is “encoded” in a representation, but to highlight the difference between information that is present (probe-recoverable) and information that is accessible (decoder-extractable).
+Rate-distortion-perception (RDP) in ML.
+Blau and Michaeli (2019) formalized perceptual quality as a distributional constraint: reconstructions are perceptually good when their distribution matches the real data distribution.
+The text-trained decoder imposes a perception constraint (text-aligned directions), forcing the adapter to compromise modality-specific information to satisfy the constraint.
+We draw on this framework in §3, where the decoder takes the role of the perceptual evaluator.
 3 Problem formulation
-Having described the phenomenon, we now formalize the architecture and the mismatch that causes it.
-Architecture. We consider the standard multimodal LLM pipeline:
-$X\xrightarrow{E_{\phi}}H\xrightarrow{P_{\theta}}Z\xrightarrow{L_{\psi}(\cdot,C)}\hat{Y}$ (1)
-where $X$ is a non-text input (speech waveform or image), $E_{\phi}$ is a pre-trained encoder (e.g., Whisper (Radford et al., 2023), CLIP (Radford et al., 2021)), $P_{\theta}$ is a learned adapter (linear or MLP projection), $Z\in\mathbb{R}^{d}$ are the hidden states consumed by the LLM $L_{\psi}$, $C$ is the text context (prompt), and $\hat{Y}$ is the generated output.
-At inference, all components are fixed: the scoring rule $q_{\psi}$ is whatever the training procedure produced.
-In the common adapter paradigm, only $P_{\theta}$ is trained while $E_{\phi}$ and $L_{\psi}$ are pre-trained; in end-to-end approaches, all components are jointly trained.
-Either way, at inference the decoder’s scoring rule is fixed, and our theory describes what that fixed scoring rule can extract.
-The specific form of $P_{\theta}$ varies (linear, MLP, Q-Former (Dai et al., 2023), perceiver, codebook), but the mismatch depends on the scoring rule, not the adapter type (§6.1).
-Text law and modal law.
-Let $q_{\psi}(y|z,c)=L_{\psi}(z,c)$ denote the LLM’s next-token distribution at inference.
+Architecture. We consider the typical multimodal LLM pipeline:
+$X\xrightarrow{E}H\xrightarrow{A}Z\xrightarrow{L(\cdot,C)}\hat{Y}$ (1)
+where $X$ is a non-text input (speech waveform or image), $E$ is a pre-trained encoder (e.g., Whisper (Radford et al., 2023), CLIP (Radford et al., 2021)), $A$ is a learned adapter, $Z\in\mathbb{R}^{d}$ are the hidden states consumed by the LLM $L$, $C$ is the text context (prompt), and $\hat{Y}$ is the generated output.
+The specific form of $A$ varies (linear, MLP, Q-Former (Dai et al., 2023), perceiver, codebook), but the mismatch depends on the scoring rule, not the adapter type (§6.1).
+Scoring rule, text law and modal law.
+Let $q(y|z,c)$ denote the LLM’s next-token distribution at inference.
+This is the fixed scoring rule resulting from training.
 During pre-training, the LLM processes hidden states drawn from the text law $P_{\!T}(C,Z,Y)$.
 When processing adapted non-text inputs, the hidden states follow a different modal law $P_{\!M}(C,Z,Y)$.
-Assumption 1 (Shared marginal).
-The text and modal laws share the same prompts and target tokens: $P_{\!M}(C,Y)=P_{\!T}(C,Y)$.
-In other words, the task is the same (e.g., “transcribe this speech”), and only the internal representations $Z$ differ.
-This isolates the effect of the representation distribution from confounds in the task itself.
-Modality collapse.
-We define modality collapse as the systematic degradation or stagnation of non-text information through a text-trained decoder, depending on the mismatch intensity.
-For an attribute $A_{\tau}$ of type $\tau$ (e.g., speaker identity, emotion), the information accessibility gap is:
-$\Delta_{\text{access}}(\tau)=I(Z;A_{\tau})-\mathrm{GMI}_{P_{\!M}}(A_{\tau}\mid\psi)$ (2)
-where $I(Z;A_{\tau})$ is the mutual information (recoverable by an optimal probe) and $\mathrm{GMI}_{P_{\!M}}(A_{\tau}|\psi)$ is the information actually extractable by the decoder’s fixed scoring rule.
-When $\Delta_{\text{access}}=0$, the decoder extracts everything the probe can; when $\Delta_{\text{access}}$ is large, information is present but inaccessible.
+We assume the text and modal laws share the same prompts and target tokens ($P_{\!M}(C,Y)=P_{\!T}(C,Y)$), so the only difference is in the representations $Z$.
+The quantitative bound (§4) additionally requires that the decoder’s log-score is Lipschitz continuous and that representations occupy a bounded region (Assumptions 23 in Appendix A.2, validated empirically in §5.6).
+Formalizing modality collapse.
+The earlier intuitive explanation of modality collapse (information is available but the decoder cannot use it) can be stated precisely.
+For an attribute $S_{\tau}$ of type $\tau$ (e.g., speaker identity, emotion), we define the information accessibility gap as:
+$\Delta_{\text{access}}(\tau)=I(Z;S_{\tau})-\mathrm{GMI}_{P_{\!M}}(S_{\tau}\mid q)$ (2)
+where $I(Z;S_{\tau})$ is the information about $S_{\tau}$ present in the representation and $\mathrm{GMI}_{P_{\!M}}(S_{\tau}|q)$ is the information actually extractable by the decoder’s fixed scoring rule.
+When $\Delta_{\text{access}}=0$, the decoder extracts everything present; when $\Delta_{\text{access}}$ is large, information is present but inaccessible.
+[…]
 Text-aligned encoders as correlation mappings.
 Contrastive encoders like CLIP (Radford et al., 2021) and SigLIP (Zhai et al., 2023) do not merge the modality manifolds (the modality gap persists (Liang et al., 2022)), but they align the informative directions: attributes that co-occur with text descriptions occupy dimensions also used by text representations.
-When such an encoder feeds an LLM, the modal law $P_{\!M}$ has high overlap with $P_{\!T}$ along informative dimensions, so the distributional distance is small and the accessibility gap is correspondingly small (§A.2 formalizes this via the Wasserstein distance).
+When such an encoder feeds an LLM, the modal law $P_{\!M}$ has high overlap with $P_{\!T}$ along informative dimensions, so the distributional distance is small and the accessibility gap is correspondingly small (we formalize this in terms of the Wasserstein distance in §A.2).
 Non-contrastive encoders like DINOv2 (Oquab et al., 2024) lack this alignment: their informative directions may be orthogonal to $P_{\!T}$, making the information effectively inaccessible to the decoder.
@@
 4 Theoretical framework
-The formulation above makes the information loss quantifiable. We establish three results (formal statements, assumptions, and proofs in Appendix A):
+The mismatched decoder framing makes three concrete predictions.
+First, since the decoder was trained on text, there should be an upper limit on how much non-text information it can extract, regardless of how much is present in the representation.
+Second, this limit should get worse as the modality representations diverge from text.
+Third, a simple classifier (a linear probe) trained directly on the representations should not be subject to the same limit, since it is not constrained by a text-trained scoring rule.
+We formalize each below (formal statements, assumptions, and proofs in Appendix A):
 1.
-The decoder has a ceiling.
-Standard mutual information $I(Z;Y)$ measures how much information $Z$ contains about $Y$ assuming an optimal decoder.
-But the LLM is not optimal for non-text inputs.
-The relevant quantity is instead the Generalized Mutual Information (GMI) (Merhav et al., 1994; Scarlett et al., 2020), which measures the maximum rate extractable by a fixed decoder with a given scoring rule.
-Regardless of how much information a probe can extract from $Z$, the decoder can never exceed $\mathrm{GMI}_{P_{\!M}}(q_{\psi})$ (Theorem 1).
+The decoder has an upper bound.
+The standard mutual information $I(Z;Y)$ represents the information $Z$ contains about $Y$ assuming an optimal decoder.
+But the LLM is not an optimal decoder for non-text inputs, given a text-focused training objective.
+Instead, the relevant quantity is the Generalized Mutual Information (GMI) (Merhav et al., 1994; Scarlett et al., 2020), the maximum rate achievable with a fixed scoring rule.
+No matter how much information any probe can get from $Z$, the decoder is always bounded by $\mathrm{GMI}_{P_{\!M}}(q)$ (Theorem 1).
 2.
 The ceiling drops with distributional distance.
-When modal representations $P_{\!M}$ differ from the text representations $P_{\!T}$ the decoder was trained on, the GMI degrades.
-We bound this degradation (Theorem 2): it scales with the product of two empirically measurable quantities: (i) the Wasserstein distance $W_{\!1}(P_{\!M},P_{\!T})$, measuring how far modal representations are from text representations, and (ii) the decoder’s Lipschitz constant $L_{\log}$, measuring how sensitive its log-probability is to shifts in the input.
-The bound makes a clear prediction: models with larger $L_{\log}\cdot W_{\!1}$ should show more degradation, which we validate in §5.5.
+The GMI decreases when the modal representations $P_{\!M}$ differ from the text representations $P_{\!T}$ the decoder was trained on.
+How much it drops depends on two measurable quantities: how far the modal representations are from text (the Wasserstein distance $W_{\!1}$), and how sensitive the decoder is to that difference (the Lipschitz constant $L_{\log}$).
+The drop is bounded by their product (Theorem 2).
+Both quantities can be estimated from a trained model ($W_{\!1}$ from the representations, $L_{\log}$ from gradient norms), so the bound yields a testable prediction: models with larger $L_{\log}\cdot W_{\!1}$ should show more degradation.
+We validate this ranking in §5.6.
+In the high-mismatch regime the bound is numerically loose, but it remains an ordinal diagnostic: it correctly ranks model–modality pairs by collapse severity and predicts which information types degrade (Appendix A.3 tightens the prefactor from ${\sim}e^{150}$ to ${\sim}e^{7}$).
 […]
+The three propositions above establish the perception-distortion face of the RDP tradeoff: as the distributional distance $W_{\!1}(P_{\!M},P_{\!T})$ grows, accessible information (GMI) decreases, and the decoder is disproportionately more affected than a probe.
 5 Experimental validation
-The bound in Theorem 2 makes three quantities empirically measurable: (i) the Lipschitz constant $L_{\log}$, estimated via gradient norms, measuring the decoder’s sensitivity to input perturbations; (ii) the Wasserstein distance $W_{\!1}(P_{\!M},P_{\!T})$, estimated between modal and text hidden states at matched layers; and (iii) the mode alignment spectrum $\tilde{\alpha}(u_{k})$, which decomposes $W_{\!1}$ geometrically by identifying which directions carry the mismatch.
-We validate each in turn.
 5.1 Setup
 Models.
-We study five multimodal LLMs spanning two modalities (Table 1).
+We investigate five multimodal LLMs across two modalities (Table 1).
 For speech: Ultravox-v0.6 (Fixie AI, 2024) (Whisper encoder, MLP adapter with SwiGLU, Llama-3.1-8B) and Qwen2-Audio-7B (Chu et al., 2024) (Whisper encoder, linear adapter, Qwen2-7B).
-For vision: LLaVA-v1.5-7B (Liu et al., 2023) (CLIP-ViT, MLP adapter, Vicuna-7B) and two Prismatic VLMs (Karamcheti et al., 2024) sharing the same MLP adapter and Vicuna-7B backbone but differing only in their vision encoder: DINOv2 (not text-aligned) and SigLIP (text-aligned via contrastive loss).
+For vision: LLaVA-v1.5-7B (Liu et al., 2023) (CLIP-ViT, MLP adapter, Vicuna-7B) and two Prismatic VLMs (Karamcheti et al., 2024) with the same MLP adapter and Vicuna-7B backbone but with different vision encoders: DINOv2 (not text-aligned) and SigLIP (text-aligned via contrastive loss).
 Model Encoder Text-aligned? LLM
@@
 Probing protocol.
-We extract representations at four hook points per model: encoder output, adapter output, LLM layer 16, and LLM final layer (after LayerNorm).
-Representations are mean-pooled across sequence positions (restricting to audio-only tokens yields nearly identical results; Appendix Table 9).
-We train $\ell_{2}$-regularized logistic regression probes (5 seeds, 80/20 stratified split, $z$-score normalization) for each combination of hook point, information type, and dataset.
+We consider four hook points per model: encoder output, adapter output, LLM layer 16, and LLM final layer (after LayerNorm).
+Representations are mean-pooled across sequence positions.
+Ultravox also provides audio token boundaries, but pooling over audio-only tokens results in nearly identical trends (see Appendix Table 9).
+For each hook point, information type and dataset combination, we train $\ell_{2}$-regularized logistic regression probes (5 seeds, 80/20 stratified split, $z$-score normalization).
 Information types.
@@
 Vision: lexical (caption-based word prediction), object category (69-way), super-category (12-way).
-The Prismatic pair (shared architecture, different encoder) and the two speech models (shared encoder, different LLM) provide controlled contrasts that isolate specific variables.
+The Prismatic pair (shared architecture, different encoder) and the two speech models (shared encoder, different LLM) serve as controls that allow us to isolate these specific variables.
+5.2 The information accessibility gap in practice
+Linear probes across layers show that for all 5 models, non-text information is not destroyed by the LLM (Appendix Tables 5 and 7): at the final layer, speaker identity is $22\times$ above chance for Ultravox and object categories are $55\times$ above chance in Prismatic-D (.792.792 vs.
+.014.014).
+The information is preserved.
+Does the decoder use it?
+The answer depends on text-alignment.
+In both speech models, the LLM approximately doubles lexical information (++92% Ultravox, ++95% Qwen2-Audio), but degrades speaker identity (-8% to -39%).
+When the encoder is text-aligned (CLIP in LLaVA, SigLIP in Prismatic-S), $P_{\!M}\approx P_{\!T}$ and all information types improve (++5–8%).
+The decoder amplifies what it was trained on and is indifferent to the rest.
+The LISTEN benchmark (Chen et al., 2026) independently observed this pattern: audio LLMs predict emotions based on what the speaker says rather than how they say it.
+Our framework explains why: the text-shaped scoring rule amplifies lexical content (++92–95%) while remaining indifferent to acoustic features, so emotion predictions are driven by word choice rather than prosody.
+The same pattern appears in vision, though in a weaker mismatch regime (Appendix Table 10).
+Three visual features not present in text (object count, object size, spatial distribution) show flat or modest improvement through the LLM ($L_{\log}\cdot W_{\!1}\approx 13$$54$ vs.
+$162$ for speech), consistent with stagnation unlike the sharp degradation in speech.
+5.3 Mode alignment
+The bound describes how much information is lost; mode alignment tells us the directions in which it is lost.
+We perform a principal component analysis of the adapter’s output representations under the modal distribution, yielding a covariance matrix $\Sigma_{M}$ with eigenvectors $u_{k}$ and eigenvalues $\lambda_{k}$, sorted by decreasing eigenvalue.
+We refer to each eigenvector as a mode: Mode 0 is the direction along which the adapter’s output varies the most, Mode 1 the next, and so on.
+For each mode, we ask: has the decoder seen variation along this direction during training?
+The alignment score
+$\rho_{k}=u_{k}^{\top}\Sigma_{T}\,u_{k}\,/\,\lambda_{k}$
+is the Rayleigh quotient of the text covariance $\Sigma_{T}$ evaluated along the modal eigenvector $u_{k}$, normalized by the modal eigenvalue $\lambda_{k}$.
+The numerator measures how much the text distribution varies along direction $u_{k}$; the denominator is how much the modality varies along the same direction.
+When $\rho_{k}=1$, the text and modal distributions have equal variance along $u_{k}$, so the decoder has been trained on exactly this kind of variation and can use it.
+When $\rho_{k}\approx 0$, the modality varies along a direction where text has almost no variance: the decoder has never seen input spread along this direction and cannot productively respond to it.
+We label modes with $\rho_{k}<0.5$ as modality-specific (MS); the threshold is a practical cutoff, and results are not sensitive to the exact value (Appendix Table 11).
+For non-aligned speech encoders, Mode 0 (the direction of largest adapter variance) is modality-specific ($\rho_{0}\leq 0.034$), accounting for 51–79% of adapter variance.
+For Prismatic-DINOv2, Mode 0 is also modality-specific ($\rho_{0}=0.034$) but carries only 8% of variance; the MS variance is instead distributed across 53 modes (71% total; Table 11).
+In both cases, the dominant source of variation in the adapter’s output is invisible to the text distribution.
+For SigLIP, even Mode 0 is text-aligned ($\rho_{0}=0.83$): the encoder has already organized its output along text-like directions.
+As representations pass through the LLM, text-aligned variance is amplified (${\sim}150\times$) while MS variance is unaffected: MS modes go from 63.6% of adapter variance to <<1% at the final layer in Ultravox, not destroyed but drowned out.
 5.4 Modality-specific structure interferes with decoding
-We directly test whether the decoder’s scoring rule functionally depends on modality-specific (MS) directions by projecting them out at inference time and measuring the change in cross-entropy.
-Unlike probing, which tests representational presence, this is a direct intervention on the decoder’s input.
-We decompose the adapter output into eigenmodes (principal directions of variance) and ask: how much of this variance does the text distribution also use?
-Each eigenmode $u_{k}$ gets a text alignment score $\tilde{\alpha}(u_{k})=u_{k}^{\top}\Sigma_{T}u_{k}/\lambda_{k}$, the fraction of its variance that text representations share (§5.6).
-Modes with $\tilde{\alpha}<0.5$ are modality-specific (MS): most of their variance comes from the non-text modality.
-We then surgically remove selected directions from the decoder’s input ($z^{\prime}=z-\sum_{k\in S}(z^{\top}u_{k})\,u_{k}$) and measure the effect on cross-entropy.
-We compare three conditions (200 samples per model): (i) ablate all MS modes, (ii) ablate a matched number of top text-aligned (TA) modes, (iii) no ablation.
+Probes show that non-text information is present; causal ablation tests whether the decoder is impacted by it.
+Unlike concept erasure methods such as LEACE (Belrose et al., 2023), which eliminate directions predictive of a labeled attribute, we ablate directions identified by distributional mismatch without concept labels.
+For each mode classified as MS ($\rho_{k}<0.5$) in §5.3, we remove its contribution from the representation: $z^{\prime}=z-\sum_{k\in S}(z^{\top}u_{k}),u_{k}$, and compare three conditions (200 samples per model): (i) ablate all MS modes, (ii) ablate a matched number of top TA modes, (iii) no ablation.
 […]
@@
 Random 11 2.8 ++0.3
+Qwen2-Audio MS (all) 5 97.6 -0.1
+TA-matched 5 1.8 ++1.0
+Random 5 0.1 ++0.2
 Prismatic-D MS (all) 53 71.0 -11.1
@@
 Table 2: Causal ablation at the adapter output.
-Eigenmodes are classified as modality-specific (MS; $\tilde{\alpha}<0.5$) or text-aligned (TA), and matched subsets are projected out.
+Eigenmodes are classified as modality-specific (MS; $\rho_{k}<0.5$) or text-aligned (TA), and matched subsets are projected out.
 Random controls ablate the same number of modes chosen uniformly at random (seed-averaged).
@@
 Random ablation gives intermediate effects proportional to variance removed.
-For SigLIP, few modes are MS and all effects are $1\%$.
-MS structure is not merely unused; its presence degrades decoding.
-Unlike concept erasure methods such as LEACE (Belrose et al., 2023), which project out directions predictive of a labeled attribute, our ablation targets directions identified by distributional mismatch between modal and text covariances, without requiring any concept labels.
-The question is not whether a concept can be erased, but whether the decoder uses these distribution-mismatched directions at all.
-Table 2 reveals a stark asymmetry across all three models.
-For Ultravox, removing 11 MS modes (63.6% of variance) slightly improves cross-entropy (-1.4%, t=7.0t=-7.0), while removing a matched TA subset has negligible effect (-0.3%, t=2.2t=-2.2).
-The Prismatic-D result is even more striking: ablating 53 MS modes (71% of variance) reduces loss by 11.1% (t=25.2t=-25.2), while ablating the top 47 TA modes (29% of variance) has no measurable effect (-0.07%, t=0.4t=-0.4, $p=0.69$).
-The decoder is functionally indifferent to the directions where most adapter variance concentrates, precisely as the mismatched-decoder theory predicts.
-For Prismatic-S (text-aligned), only 14 of 100 modes are MS (24.6% of variance) and all ablation effects are below 1%, consistent with SigLIP producing representations the decoder already expects.
+For SigLIP, few modes are MS and all effects are <1{<}1%.
+[…]
+For Prismatic-S (text-aligned), only 14 of 100 modes are MS (24.6% of variance) and all ablation effects are below 1%, consistent with SigLIP representations already matching what the decoder expects.
 5.5 Controlled experiment: text-aligned vs. non-aligned encoders
-The causal ablation establishes that MS directions interfere with decoding.
-We now isolate why some models have more MS variance than others, using the Prismatic pair as a controlled comparison that holds everything constant except encoder text alignment.
-Both Prismatic VLMs share the same MLP adapter, Vicuna-7B backbone, and training recipe.
-The only difference is the vision encoder: DINOv2 (Oquab et al., 2024) (no text alignment) versus SigLIP (Zhai et al., 2023) (contrastive text–image loss).
-If our theory is correct, SigLIP should produce representations closer to $P_{\!T}$, yielding smaller $W_{\!1}(P_{\!M},P_{\!T})$ and less information loss.
-The results confirm the prediction (Appendix Figure 1, Table 10).
-With DINOv2, visual information passes through the LLM essentially unchanged (-0.0% object category).
-With SigLIP, the same information types improve (++1.9% object category, ++2.3% super-category).
-This pattern replicates across modalities: in speech, Whisper-based encoders show dramatic lexical recovery (++92–95%) alongside speaker collapse (-8% to -39%); in LLaVA (CLIP encoder, text-aligned), all information types improve (Appendix Table 7).
-5.4 The information accessibility gap in practice
-The causal ablation (§5.2) establishes that the decoder does not use MS directions, and the controlled comparison (§5.3) traces this to encoder text alignment.
-Probes confirm the information has not been destroyed (Appendix Tables 75).
+The causal ablation shows that MS directions influence decoding.
+Now, we explain why some models have more MS variability than others.
+The Prismatic pair is ideal for this, as it controls for all factors except the text-alignment of the encoder.
+Both Prismatic VLMs use the same MLP adapter, Vicuna-7B backbone architecture and training protocol.
+The difference lies in the vision encoder: DINOv2 (Oquab et al., 2024) (no text alignment) versus SigLIP (Zhai et al., 2023) (contrastive text–image loss).
+According to our hypothesis, the representations learned by SigLIP should be more similar to $P_{\!T}$, leading to smaller $W_{\!1}(P_{\!M},P_{\!T})$ and less information loss.
+This is what we see (Appendix Table 10): the split is sharper for non-textual attributes (object count, object size, spatial spread), which cannot be inferred from captions alone.
+For these features, SigLIP shows consistent improvement through the LLM (+4.5–8.2%) while DINOv2 gains less (+1.3–4.7%), with the largest gap on object count (+8.2% vs.
++3.4%).
 […]
-Three non-textual visual attributes (object count, size, spatial spread) derived from COCO annotations but absent from captions confirm the pattern in a milder mismatch regime ($L_{\log}\cdot W_{\!1}\approx 13$$54$ vs.
-$162$ for speech), explaining stagnation rather than degradation (Appendix Table 10).
 5.6 Validating the bound
-Theorem 2 predicts that GMI degradation scales with $L_{\log}\cdot W_{\!1}$.
-We estimate both quantities empirically (full results in Appendix Table 8).
-$L_{\log}$ is finite for all nine model-dataset combinations, confirming the Lipschitz log-score assumption (Assumption 2, Appendix A.2).
-Three patterns emerge: (i) $L_{\log}$ is content-dependent, decreasing as audio moves away from text ($9.08$ speech $\to$ $3.66$ environmental sounds in Ultravox); (ii) $L_{\log}$ varies $30\times$ across architectures ($0.29$ LLaVA to $9.08$ Ultravox); (iii) Prismatic-S has $2.8\times$ higher $L_{\log}$ than Prismatic-D ($1.12$ vs.
-$0.40$), consistent with the LLM being more responsive to text-aligned representations.
-$W_{\!1}$ ranges from 17.8 to 50.4 across models.
+$L_{\log}$ is finite for all nine model-dataset pairs, validating the Lipschitz assumption (Appendix Table 8).
+It varies with content ($9.08$ for speech to $3.66$ for environmental sounds in Ultravox) and differs $30\times$ across architectures ($0.29$ LLaVA to $9.08$ Ultravox).
+Prismatic-S has $2.8\times$ larger $L_{\log}$ than Prismatic-D ($1.12$ vs.
+$0.40$): the decoder is more responsive to representations that resemble text.
 Gradient isotropy.
-The mean gradient norm along modality-specific directions ($g_{\text{MS}}$) is nearly identical to that along text-aligned directions ($g_{\text{TA}}$): $g_{\text{TA}}/g_{\text{MS}}=0.94\times$, Spearman $r=0.11$, $p=0.28$.
-This validates the scalar $L_{\log}$ assumption: the decoder responds with similar magnitude to perturbations in any direction.
-But MS modes carry structured variance that pushes predictions away from text distributions, while TA modes carry variance the decoder was trained to exploit.
-Same sensitivity, different semantic effect.
-The product $L_{\log}\cdot W_{\!1}$ is largest for Ultravox ($162$), which shows the most dramatic degradation (-39% speaker), and smallest for LLaVA ($13.4$), where all information types improve.
-Prismatic-S has a higher product ($53.7$) than Prismatic-D ($20.2$) despite showing less degradation; this reflects higher decoder responsiveness to text-aligned representations rather than greater mismatch.
-The bound’s prefactor is also smaller for Prismatic-S because its effective support diameter is smaller ($D=16.2$ vs.
-$35.7$).
-The raw bound uses the ambient representation diameter $D$, which makes the prefactor vacuously large.
-The support-restricted bound (Appendix A.3) allows replacing $D$ with the effective support diameter $D_{\text{eff}}$, estimated via the participation ratio of the pooled representation covariance.
-This yields $L_{\log}\cdot D_{\text{eff}}\in[3,7]$, making the prefactor moderate ($\sim$20–1100).
-The bound is structural rather than tight: it identifies $L_{\log}\cdot W_{\!1}$ as the governing quantity and correctly predicts which models collapse most.
-Per-type prediction.
-The theory correctly predicts the aggregate pattern: text-aligned information recovers or improves, while non-text types degrade or stagnate.
-A finer prediction (degradation correlating with each type’s MS-variance share) does not hold at this granularity (Spearman $\rho=-0.40$, $p=0.60$), likely because the content-dependent $L_{\log}$ (which varies $2.5\times$ across audio types) modulates degradation beyond geometric alignment alone.
-5.6 Mode alignment
-The bound explains how much information is lost; mode alignment explains which directions carry the mismatch.
-We decompose the adapter output into its principal directions of variance (eigenmodes of $\Sigma_{M}$) and measure, for each direction, the fraction of its variance that the text distribution shares: $\tilde{\alpha}(u_{k})=u_{k}^{\top}\Sigma_{T}u_{k}/\lambda_{k}$.
-A score near 1 means the direction is shared with text; near 0 means it is unique to the modality (Appendix Table 11, Figure 2).
-For non-aligned encoders, the largest direction (Mode 0) is modality-specific ($\tilde{\alpha}\leq 0.034$), capturing 51–79% of adapter variance.
-Most of what the adapter transmits is invisible to the text distribution.
-For SigLIP, even Mode 0 is shared with text ($\tilde{\alpha}=0.83$).
-As representations pass through the LLM, text-aligned variance is amplified (${\sim}150\times$) while MS variance remains flat: MS modes drop from 63.6% of adapter variance to <<1% at the final layer in Ultravox, not destroyed but drowned out.
+To check whether the decoder selectively ignores MS directions at the gradient level, we compute the norm of the decoder’s loss gradient when the input representation is perturbed along each eigenmode direction.
+Let $g_{k}=|\nabla_{u_{k}}\mathcal{L}|$ be the gradient norm for mode $k$.
+The mean gradient norm for TA modes ($\bar{g}_{\text{TA}}$) and MS modes ($\bar{g}_{\text{MS}}$) are nearly equal ($\bar{g}_{\text{TA}}/\bar{g}_{\text{MS}}=0.94$, Spearman $r=0.11$, $p=0.28$): the decoder does not selectively ignore MS directions.
+This confirms the scalar $L_{\log}$ assumption.
+The product $L_{\log}\cdot W_{\!1}$ ranks models by collapse severity.
+Ultravox ($L_{\log}\cdot W_{\!1}=162$) shows the most degradation (-39% speaker); LLaVA ($13.4$) shows none.
+Prismatic-S has a higher product ($53.7$) than Prismatic-D ($20.2$) despite less degradation, reflecting greater decoder responsiveness to text-aligned representations rather than greater mismatch.
+As noted in §4, the bound is an ordinal diagnostic: it correctly ranks models and predicts which information types degrade, even in the high-mismatch regime where it is numerically loose (Appendix A.3A.4).
 6 Discussion
@@
 The fix is objective-side, not encoder-side.
-If modality collapse is a consequence of the scoring rule being text-shaped, the natural intervention is to reshape the scoring rule.
-Low-rank adaptation (LoRA; Hu et al., 2022) on the LLM adjusts $q_{\psi}$ to better match $P_{\!M}$, effectively reducing the functional mismatch.
-Preliminary evidence from concurrent work (Billa, 2026) is consistent: applying LoRA to the LLM backbone reduces text dominance by 23.9%, while adapter-only training increases text dominance by 26.5%.
-This is predicted by Theorem 2: LoRA reduces the functional mismatch between $P_{\!M}$ and $P_{\!T}$ by reshaping the scoring rule, whereas adapter-only training leaves the decoder’s text-shaped sensitivity unchanged.
+If modality collapse is a result of the scoring rule being text-shaped, a logical solution would be to reshape the scoring rule.
+Low-rank adaptation (LoRA; Hu et al., 2022) on the LLM modifies $q$ to make it closer to $P_{\!M}$, effectively reducing the functional mismatch.
+Preliminary evidence from concurrent work on audio-text conflict supports this: applying LoRA to the LLM backbone decreases text dominance, whereas adapter-only training increases it.
+[^1] Citation omitted for double-blind review.
+This is predicted by Theorem 2: LoRA reduces the functional mismatch between $P_{\!M}$ and $P_{\!T}$ by reshaping the scoring rule, whereas adapter-only training does not change the decoder’s text-shaped sensitivity.
 However, LoRA on the decoder is necessary but not sufficient: the training objective must explicitly target non-text information.
-We demonstrate this with two contrasting experiments on Ultravox, using the same LoRA configuration (r=16r{=}16, $\alpha{=}32$, targeting q/k/v/o_proj).
-Negative example (text-centric objective).
-We apply a LoRA checkpoint trained for audio-text conflict resolution (ALME; Billa, 2026) and measure linear probe accuracy on LibriSpeech.
-Speaker identity accuracy is unchanged (55.0% vs.
-54.5% base, within noise), as is lexical accuracy (53.5% vs.
-52.6%).
-The ALME objective rewards resolving text-audio conflicts, not preserving speaker or acoustic attributes, so its gradient signal never reaches modality-specific directions.
-Positive example (emotion objective).
-We train a LoRA on the LLM backbone with a forced-choice emotion detection objective on CREMA-D (6 emotions, 7,442 clips, standard causal LM loss on assistant tokens only).
-Generation accuracy on held-out samples improves from 17.3% to 61.8% (Table 3).
-Critically, the emotion probe at llm-final improves from .557.557 to .632.632 (++7.5%), while the speaker probe remains at chance (.135.135 vs.
-.137.137) and lexical accuracy remains near-ceiling (.994.994).
-Upstream layers (encoder, adapter) are identical across conditions, confirming the change is entirely in the LLM.
-The LoRA reshapes the scoring rule so that, at inference, the decoder responds to emotion-relevant directions while leaving others untouched, as Theorem 2 predicts: the emotion objective reduces $W_{\!1}$ along emotion-correlated dimensions without affecting orthogonal ones.
+We show this with a LoRA experiment on Ultravox (r=16r{=}16, $\alpha{=}32$, targeting q/k/v/o_proj).
+LoRA with emotion objective.
+We fine-tune with LoRA on the LLM backbone using a forced-choice emotion detection objective on CREMA-D (6 emotions, 7,442 clips, standard causal LM loss on assistant tokens only).
+We evaluate the effect of LoRA on two measures: linear probe accuracy at the LLM’s final layer (measuring what information is accessible in the representation) and task accuracy (measuring whether the decoder actually uses it in its output).
+Task accuracy on held-out samples improves from 17.3% to 61.8% (Table 3).
+The emotion probe at llm-final improves from $557\pm.006$ to $632\pm.011$ (++7.5%), while the speaker probe remains at chance ($135\pm.006$ vs.
+$137\pm.008$) and lexical accuracy stays near-ceiling ($994\pm.001$).
+Upstream layers (encoder, adapter) are shared between conditions, so this is purely a change to the LLM.
+LoRA modifies the scoring rule such that at inference, the decoder will respond to the subset of directions that are emotion-relevant and leave the rest unchanged.
+Consistent with Theorem 2, the emotion objective reduces $W_{\!1}$ along the subset of directions correlated with emotion, while leaving all others unchanged.
 Table 3: LoRA intervention on Ultravox (CREMA-D, llm-final probe accuracy).
 The emotion LoRA selectively improves emotion accessibility without meaningfully affecting speaker or lexical probes.
-Generation accuracy measures forced-choice emotion detection on 1,002 held-out samples.
+Task accuracy measures forced-choice emotion detection on 1,002 held-out samples.
 The encoder and adapter outputs are identical across conditions (LoRA modifies only the LLM).
-Condition Emotion Speaker Lexical Gen. acc.
-Base Ultravox .557 .137 1.000 17.3%
+Condition Emotion (probe) Speaker (probe) Lexical (probe) Task acc.
+Base Ultravox .557 .137 .999 17.3%
 + Emotion LoRA .632 .135 .994 61.8%
-$\Delta$ ++7.5% -0.2% -0.6% ++44.5%
-Together, the negative and positive examples confirm that the training objective determines which dimensions of the scoring rule are reshaped.
-A text-centric objective (ALME) leaves the decoder blind to non-text directions; an objective targeting a specific non-text attribute (emotion) teaches the decoder to respond to exactly that dimension.
+$\Delta$ ++7.5% -0.2% -0.5% ++44.5%
 Text-aligned encoders are a workaround, not a solution.
-Contrastive encoders like CLIP and SigLIP achieve low mismatch by mapping inputs onto text-correlated dimensions, but this is a projection, not a preservation: modality-specific information with no textual correlate is discarded at the encoder.
-The Prismatic comparison confirms this: SigLIP’s visual information “improves” through the LLM only because it was pre-projected onto text-like directions.
-CLIP gives the LLM access to textual correlates of visual information, not visual information itself.
+Contrastive encoders like CLIP and SigLIP achieve low mismatch because they project the input data into the space of text-correlated features.
+The alignment is effective but selective in that it preserves what co-occurs with text and discards the rest.
+The Prismatic comparison shows that SigLIP’s visual representations “enhanced” by the LLM in fact only looked better because they were pre-projected into text-like directions.
+CLIP provides the LLM with visual features organized along text-correlated dimensions; genuinely visual information (texture, spatial layout) is discarded upstream.
 Models must be explicitly trained to use non-text modalities.
-A text-trained decoder responds only to directions it was trained on (§5.6).
+A text-only decoder will not listen to signals it was not trained on (§5.3).
 No adapter can make it sensitive to directions it has never seen.
-To build models that genuinely leverage modality-specific information, not just its textual shadow, the training objective must reward extracting non-text directions, whether through LoRA, full multimodal pre-training, or other objective-side interventions.
-Without this, non-textual information remains hidden in plain sight.
+If we want a model to use modality-specific knowledge, rather than merely its text representation, we need to provide a training signal for it to do so, whether through LoRA, full multimodal pretraining, or some other objective-side intervention such as a modality-specific loss.
+Otherwise the information is just not available.
 The framework is architecture-agnostic.
-Our experiments use models with explicit adapter modules, but the analysis does not depend on them.
-The bound is a property of the decoder’s scoring rule, not of any particular architectural component.
-Any system where a text-trained decoder processes non-text representations faces the same constraint, whether the representations arrive through a linear projection, an MLP, a Q-Former, a discrete codebook (Borsos et al., 2023; Défossez et al., 2024), or no explicit adapter at all.
-The LLM’s own layers implicitly serve as the adapter.
-Discrete-token approaches add a quantization bottleneck that can only increase $W_{\!1}(P_{\!M},P_{\!T})$, making our bound conservative.
-The core mechanism is the mismatch between the decoder’s training distribution and its inputs at inference, and that mechanism is universal.
+The bound we present depends on the decoder’s scoring rule not the model’s architecture.
+The source of representations, a linear projection, an MLP, a Q-Former, a discrete codebook (Borsos et al., 2023; Défossez et al., 2024), or no explicit adapter at all (i.e., the LLM’s own layers act as the adapter), is immaterial.
+Discrete token methods introduce an additional quantization bottleneck that can only increase $W_{\!1}(P_{\!M},P_{\!T})$, making our bound more conservative.
+The essential phenomenon is the distributional shift the decoder sees between training and inference.
 6.2 Limitations
 Probes as information measures.
-Linear probes are necessary but not sufficient (Belinkov, 2022), but the key asymmetry (more information is linearly recoverable than the decoder uses) holds regardless.
-We estimate $L_{\log}$ via p95 gradient norms, not the true supremum.
-Additional limitations (scope, Assumption 1) are in Appendix D.
+Linear probes provide a necessary but not sufficient condition for information presence (Belinkov, 2022), but the main asymmetry (more information can be linearly recovered than the decoder uses) holds either way.
+We can only approximate $L_{\log}$ via p95 gradient norms, not compute the true supremum.
+Other caveats (in scope, Assumption 1) can be found in Appendix C.
 […]
+Scale.
+All five models studied are in the 7–8B parameter range.
+The bound (Theorem 2) depends on the scoring rule’s sensitivity ($L_{\log}$) and the distributional distance ($W_{\!1}$), not on parameter count, so the theoretical constraint is scale-invariant.
+For larger models the bound should remain consistent with our results insofar as a text-majority-training-objective paradigm remains.
+Validating the bound at the 70B scale is a direction for future work.
 7 Conclusion
-[…]
-Building multimodal models that truly use what they perceive requires reshaping the decoder’s scoring rule.
-Text-aligned encoders appear to help but work by discarding non-textual information at the encoder.
-Our LoRA experiment shows the real fix: training with an emotion objective teaches the decoder to respond to emotion-relevant directions (++7.5% probe, ++44.5% generation accuracy) while leaving others untouched.
-The framework extends beyond the models and modalities studied here.
-The bound is a property of the decoder’s scoring rule, not of any architectural component; no explicit adapter is required.
-At inference, every trained model has a fixed scoring rule.
-If that scoring rule was shaped predominantly by text, the constraint applies regardless of architecture or training recipe.
-The question is not whether modality-specific information can be encoded (it already is) but whether the training objective teaches the decoder to use it.
-Until it does, that information remains hidden in plain sight.
+Modality collapse is caused by a failure of decoding, not encoding, as our experiments demonstrate.
+The GMI-Wasserstein bound shows theoretically that only information in the text-aligned directions is accessible to the decoder, and the rate at which information is lost depends on distributional divergence and decoder sensitivity.
+To train multimodal models that properly exploit the information they receive, it is necessary to adjust the scoring function of the decoder.
+Text-aligned encoders do appear to improve performance, but they do so by ignoring all non-text modalities at the level of the encoder.
+Our LoRA experiment confirms this: an emotion-based objective trains the decoder to be sensitive to emotion-relevant messages (++7.5% probe, ++44.5% task accuracy), without affecting irrelevant messages.
+The question, for any multimodal model, is not whether modality-specific details can be represented (they are) but whether the training objective instructs the decoder to rely on them.
 Reproducibility statement
@@
 All experiments run on a single NVIDIA RTX 3090 Ti (24GB).
+Ethics & AI disclosure
+This work analyses existing publicly available models and datasets; no new data was collected and no human subjects were involved.
+The author used Claude/Claude Code during preparation for manuscript critique, narrative feedback, literature search, and experiment implementation and debugging.
+All research design, theoretical development, experimental execution, analysis, and writing are the author’s own.
+The author takes full responsibility for all content.
 Appendix A Theoretical framework: detailed treatment
-This appendix provides the full formal development summarized in §4.
-We aim to make this self-contained: each result is preceded by an intuitive explanation, followed by the formal statement and a detailed proof with annotations.
+This appendix contains the detailed formal development outlined in §4. Each result is preceded by an informal intuitive statement followed by a complete proof.
 A.1 Generalized Mutual Information (GMI)
-The core question.
-A representation $Z$ may contain information about a target $Y$ (e.g., the next token).
-But how much of that information can a specific decoder actually use?
-Standard mutual information $I(Z;Y)$ answers this for an ideal decoder that is perfectly tuned to the data it receives.
-A text-trained LLM is not ideal for non-text inputs: it was tuned on text representations, not speech or image representations.
-We need a measure that accounts for this mismatch.
-That measure is the Generalized Mutual Information (GMI) (Merhav et al., 1994; Scarlett et al., 2020).
-We use the same notation as §3: $q_{\psi}(y|z,c)$ is the decoder’s next-token distribution (Eq.
+Overview.
+A representation $Z$ might encode some knowledge about a target $Y$ (e.g., the next token), but how much of this knowledge can a given decoder use?
+Standard mutual information $I(Z;Y)$ answers this for a decoder that is optimally matched to the representations it gets.
+A text-trained LLM is not optimally matched to non-text representations.
+It was matched to text representations, not speech representations, or image representations.
+The appropriate measure is the Generalized Mutual Information (GMI) (Merhav et al., 1994; Scarlett et al., 2020).
+We reuse the notation from §3: $q(y|z,c)$ is the decoder’s next-token distribution (Eq.
 1), $Z$ is the representation consumed by the LLM, $C$ is the text context, and $Y$ is the target token.
-Let $\ell_{\psi}(c,z,y):=\log q_{\psi}(y|z,c)$ be the decoder’s log-score (the log-probability it assigns to token $y$ given representation $z$ and context $c$), clipped at a floor $\eta=1/|\mathcal{V}|$ (where $\mathcal{V}$ is the decoder’s output vocabulary) to ensure boundedness.
-Intuition. The GMI measures how well the decoder can tell the “right” representation from random alternatives. In words:
+Let $\ell(c,z,y):=\log q(y|z,c)$ be the decoder’s log-score (the log-probability it assigns to token $y$ given representation $z$ and context $c$), clipped at a floor $\eta=1/|\mathcal{V}|$ (where $\mathcal{V}$ is the decoder’s output vocabulary) to ensure boundedness.
+Intuition. The GMI tells us how well the decoder can distinguish the “correct” representation from random alternatives:
 […]
-When the decoder reliably scores the correct representation higher than random draws, the GMI is large.
-When it treats all representations interchangeably (i.e., it cannot tell which $Z$ goes with which $Y$), the GMI collapses to zero.
-This is closely related to the Donsker-Varadhan variational representation of KL divergence (Donsker & Varadhan, 1983), applied to the decoder’s scoring function rather than an arbitrary test function.
-Formally, for a joint distribution $P$ over $(C,Z,Y)$ (either the text distribution $P_{\!T}$ or the modal distribution $P_{\!M}$ from §3), the GMI is:
-$\mathrm{GMI}_{P}(q_{\psi})=\mathbb{E}_{(C,Z,Y)\sim P}\!\left[\ell_{\psi}(C,Z,Y)-\log\mathbb{E}_{Z^{\prime}\sim P(\cdot|C,Y)}\!\left[e^{\ell_{\psi}(C,Z^{\prime},Y)}\right]\right].$ (3)
+If the decoder consistently assigns a higher score to the true representation than to a uniformly sampled one, then the GMI is large.
+If the decoder cannot distinguish which $Z$ generated which $Y$ (assigning the same score to both), then the GMI is zero.
+This is similar to the Donsker-Varadhan variational representation of KL divergence (Donsker and Varadhan, 1983), but applied to the decoder’s scoring function instead of an arbitrary test function.
+Formally, for a joint distribution $P$ over $(C,Z,Y)$ (which is either the text distribution $P_{\!T}$ or the modal distribution $P_{\!M}$ from §3), the GMI is given by:
+$\mathrm{GMI}_{P}(q)=\mathbb{E}_{(C,Z,Y)\sim P}\!\left[\ell(C,Z,Y)-\log\mathbb{E}_{Z^{\prime}\sim P(\cdot|C,Y)}\!\left[e^{\ell(C,Z^{\prime},Y)}\right]\right].$ (3)
 Theorem 1 (Accessible rate = GMI; following Scarlett et al., 2020).
-Under i.i.d.
-sampling from $P_{\!M}$ and standard measurability conditions, the maximum rate extractable by the fixed decoder $q_{\psi}$ is $R_{\mathrm{acc}}(P_{\!M},q_{\psi})=\mathrm{GMI}_{P_{\!M}}(q_{\psi})$.
-This is a direct application of the random-coding argument of Scarlett et al.
-(2020), Chapters 2–3, with constant-composition codebooks under the GMI decoder metric.
-The key insight is that the decoder $q_{\psi}$ defines a fixed scoring rule (fixed at inference, regardless of how it was trained), and the maximum rate extractable under this rule, averaged over random codebooks drawn from $P_{\!M}$, converges to $\mathrm{GMI}_{P_{\!M}}(q_{\psi})$.
-The implication is sharp: regardless of how much information a probe can extract from $Z$, the decoder can never exceed $\mathrm{GMI}_{P_{\!M}}(q_{\psi})$.
+Under i.i.d. sampling from $P_{\!M}$ and standard measurability conditions, the maximum rate extractable by the fixed decoder $q$ is $R_{\mathrm{acc}}(P_{\!M},q)=\mathrm{GMI}_{P_{\!M}}(q)$.
+This is a classical result from the mismatched decoding literature; we state it here for completeness and to establish notation.
+The proof follows directly from the random-coding argument of Scarlett et al.
+(2020), Chapters 2–3, with constant-composition codebooks and the GMI decoder metric.
+The main idea is that the decoder $q$ defines a fixed scoring rule (fixed at inference, regardless of how it was trained), and the maximum rate extractable under this rule, averaged over random codebooks distributed according to $P_{\!M}$, approaches $\mathrm{GMI}_{P_{\!M}}(q)$.
 A.2 GMI-Wasserstein bound
-We now bound the GMI degradation when the decoder sees modal representations $P_{\!M}$ instead of the text representations $P_{\!T}$ it was trained on.
-The intuition is simple: if modal representations are close to text representations, the decoder should perform nearly as well on them.
-Two properties of the decoder determine how “close” is close enough: how sensitive it is to small shifts in its input, and how large the space of possible representations is.
+Next we bound the degradation of GMI when the decoder operates on the modal representations $P_{\!M}$ rather than the text representations $P_{\!T}$ it was trained on.
+The basic idea is as follows: if modal representations are sufficiently close to text representations, then the decoder should achieve near the same performance on them.
+How “sufficiently close” is determined by two properties of the decoder: how sensitive the decoder is to small perturbations of its inputs, and how large the space of representations is.
+Theorem 2 relies on three assumptions: the shared marginal (Assumption 1), the Lipschitz condition on the decoder (Assumption 2), and bounded support (Assumption 3).
+Assumption 1 (Shared marginal).
+The text and modal laws have the same prompts and target tokens: $P_{\!M}(C,Y)=P_{\!T}(C,Y)$.
 Assumption 2 (Local Lipschitz log-score).
-There exists $L_{\log}>0$ such that for all $(c,y)$ and all $z,z^{\prime}$ in the typical-state region $\mathcal{Z}$: $|\ell_{\psi}(c,z,y)-\ell_{\psi}(c,z^{\prime},y)|\leq L_{\log}\|z-z^{\prime}\|$.
-That is, if you nudge a representation by a small amount, the decoder’s log-probability changes by at most $L_{\log}$ times that amount.
-The constant $L_{\log}$ captures how sensitive the decoder is to shifts in its input.
-Neural networks with smooth activations satisfy this locally; we estimate $L_{\log}$ empirically via gradient norms (§5.5).
+There exists $L_{\log}>0$ such that for all $(c,y)$ and all $z,z^{\prime}$ in the typical-state region $\mathcal{Z}$: $|\ell(c,z,y)-\ell(c,z^{\prime},y)|\leq L_{\log}\|z-z^{\prime}\|$.
+That is, perturbing the representation by any amount changes the log-probability assigned by the decoder by at most $L_{\log}$ times that amount.
+$L_{\log}$ encodes the sensitivity of the decoder to changes to the input.
+Neural networks with smooth activations meet this condition locally; we measure $L_{\log}$ empirically using gradient norms (§5.6).
 Assumption 3 (Bounded typical-state region).
 The typical-state region $\mathcal{Z}$ has diameter $D=\sup_{z,z^{\prime}\in\mathcal{Z}}\|z-z^{\prime}\|$, which is finite and $O(\sqrt{d})$ after LayerNorm.
-This says the representations that actually occur in practice live in a bounded region, not scattered arbitrarily through $\mathbb{R}^{d}$.
-LayerNorm ensures this by projecting all hidden states onto a hypersphere of radius $\sqrt{d}$.
+This means the representations that are actually used in practice are confined to a bounded region, rather than being distributed throughout $\mathbb{R}^{d}$.
+This is precisely what LayerNorm accomplishes: all hidden states are projected onto a hypersphere of radius $\sqrt{d}$.
 Theorem 2 (GMI-Wasserstein bound).
 Under Assumptions 1, 2, and 3:
-$|\mathrm{GMI}_{P_{\!M}}(q_{\psi})-\mathrm{GMI}_{P_{\!T}}(q_{\psi})|\leq\big(1+e^{L_{\log}\cdot D}\big)\cdot L_{\log}\cdot\mathbb{E}_{(C,Y)}\!\left[W_{\!1}\!\big(P_{\!M}(Z|C,Y),\,P_{\!T}(Z|C,Y)\big)\right].$ (4)
-Reading the bound. In words:
+$|\mathrm{GMI}_{P_{\!M}}(q)-\mathrm{GMI}_{P_{\!T}}(q)|\leq\big(1+e^{L_{\log}\cdot D}\big)\cdot L_{\log}\cdot\mathbb{E}_{(C,Y)}\!\left[W_{\!1}\!\big(P_{\!M}(Z|C,Y),\,P_{\!T}(Z|C,Y)\big)\right].$ (4)
+Intuition.
 […]
-The gap between what the decoder could extract from text representations and what it can extract from modal representations is controlled by how far apart those distributions are, amplified by how sensitive the decoder is.
-Crucially, $W_{\!1}$ measures distributional distance, not just the distance between means: aligning the average modal representation with the average text representation is not enough.
-If the modal distribution has different shape, spread, or tails, the exponential penalty factor $(1+e^{L_{\log}D})$ catches them.
-This is why text-aligned encoders reduce the bound (they project onto text-like directions, reducing both the distance and the shape mismatch) while simple mean-centering would not.
-The prefactor $(1+e^{L_{\log}D})$ arises from the “competition term” in the proof (see below).
-In the ambient space this prefactor is vacuously large ($L_{\log}\cdot D$ ranges from 10 to 150); the support-restricted bound (§A.3) tightens it using the effective support diameter.
+The difference between how much information the decoder can get from text representations and how much it can get from modal representations is that distance between distributions, multiplied by the strength of the decoder.
+But we have $W_{\!1}$ of the distributions, not the distance between the means; getting the mean of the modal distribution to the mean of the text distribution is not enough.
+If the modal distribution has a different shape, or spread, or tails, then the exponential penalty factor $(1+e^{L_{\log}D})$ gets them.
+This is why the text-aligned encoders help to bound (they project on text-like directions, so they decrease both the distance and the shape mismatch), but mean centering would not.
+The prefactor $(1+e^{L_{\log}D})$ comes from the “competition term” in the proof (below).
+In the ambient space this prefactor is trivially large ($L_{\log}\cdot D$ ranges from 10 to 150) and the support-restricted bound (§A.3) refines it in terms of the effective support diameter.
 A.2.1 Proof of Theorem 2
 Proof overview.
-The GMI has two pieces: (A) how well the decoder scores the correct representation, and (B) how well it scores random alternatives.
-We need to show that switching from text to modal representations does not change either piece by too much.
-For piece (A), the argument is simple: if the decoder’s scores change smoothly (Lipschitz assumption), then moving representations a small distance changes expected scores by a small amount.
-For piece (B), the argument is harder because piece (B) involves an exponential of the scores, which amplifies small changes.
-The bounded-diameter assumption (representations live in a finite region) keeps this amplification under control.
-Setup. Write $\ell(z)\equiv\ell_{\psi}(c,z,y)$ for fixed $(c,y)$. The GMI functional decomposes as $\Gamma(P)=A(P)-B(P)$ where:
+There are two terms in the GMI: (A) the expected score of the true representation, and (B) the expected score of a random representation.
+We must prove that the quantity in both (A) and (B) change very little when we replace the text representations with the modal representations.
+For (A), this is a corollary of the Lipschitz assumption.
+For (B), we have a potential complication of an exponential of the scores.
+However, the bounded-diameter assumption ensures that representations are all contained in a bounded region, which prevents large changes due to the exponential.
+Setup. We denote $\ell(z)\equiv\ell(c,z,y)$ for fixed $(c,y)$. The GMI functional separates as $\Gamma(P)=A(P)-B(P)$ where:
 $A(P)$ $=\mathbb{E}_{Z\sim P}[\ell(Z)]$ (direct term: expected score of the correct representation) (5)
 $B(P)$ $=\log\mathbb{E}_{Z\sim P}[e^{\ell(Z)}]$ (competition term: how well random alternatives score) (6)
-We bound $|A(P_{\!M})-A(P_{\!T})|$ and $|B(P_{\!M})-B(P_{\!T})|$ separately.
+We derive upper bounds for $|A(P_{\!M})-A(P_{\!T})|$ and $|B(P_{\!M})-B(P_{\!T})|$ separately.
 Step 1: Direct term (how much does the correct representation’s score change?).
-We want to bound how much the expected score changes when we switch from $P_{\!T}$ to $P_{\!M}$.
-The Kantorovich–Rubinstein theorem (Kantorovich & Rubinstein, 1958; Villani, 2009) says: if a function changes by at most $L$ when you move its input by 1 unit ($L$-Lipschitz), then the difference in its expected value under two distributions is at most $L$ times the Wasserstein distance between those distributions.
+We would like to bound how much the expected score changes when we go from $P_{\!T}$ to $P_{\!M}$.
+The Kantorovich–Rubinstein theorem (Kantorovich and Rubinstein, 1958; Villani, 2009) states that: if a function changes by at most L when you change its input by 1 unit ($L$-Lipschitz), then the difference in its expectation w.r.t.
+two distributions is at most L times the Wasserstein distance between the two distributions.
 Since $\ell$ is $L_{\log}$-Lipschitz (Assumption 2):
 $|A(P_{\!M})-A(P_{\!T})|=|\mathbb{E}_{P_{\!M}}[\ell]-\mathbb{E}_{P_{\!T}}[\ell]|\leq L_{\log}\cdot W_{\!1}(P_{\!M},P_{\!T}).$ (7)
-In words: if modal and text representations are close (small $W_{\!1}$) and the decoder’s scores change smoothly (small $L_{\log}$), then the expected score barely changes.
+If modal and text representations are similar (small $W_{\!1}$), and the decoder scores are smooth (small $L_{\log}$), then the expected score changes by a tiny amount.
 Step 2: Competition term (how much does the random-alternative score change?).
-This is harder because $B(P)=\log\mathbb{E}[e^{\ell(Z)}]$ involves an exponential of the score, which amplifies small differences.
-A coupling $\pi$ is a way of pairing up samples from $P_{\!M}$ and $P_{\!T}$ (imagine lining up modal representations next to the text representations they most resemble).
-For each pair $(Z_{M},Z_{T})$, define $\Delta\ell=\ell(Z_{M})-\ell(Z_{T})$: the difference in score between the paired representations.
+This is harder because $B(P)=\log\mathbb{E}[e^{\ell(Z)}]$ involves an exponential of the score, which means small differences are magnified.
+We can use a coupling $\pi$: a way of pairing up samples from $P_{\!M}$ and $P_{\!T}$ (think of arranging modal representations alongside the text representations they resemble the most).
+For each pair $(Z_{M},Z_{T})$, define $\Delta\ell=\ell(Z_{M})-\ell(Z_{T})$ as the difference in score between paired representations.
 By the Lipschitz assumption, $|\Delta\ell|\leq L_{\log}\cdot\|Z_{M}-Z_{T}\|$, and by the bounded diameter (Assumption 3), $|\Delta\ell|\leq L_{\log}\cdot D$.
@@
 $\frac{m(P_{\!M})}{m(P_{\!T})}$ $=\mathbb{E}_{\pi}\!\left[\frac{e^{\ell(Z_{M})}}{m(P_{\!T})}\right]=\mathbb{E}_{\pi}\!\left[\frac{e^{\ell(Z_{T})}}{m(P_{\!T})}\cdot e^{\Delta\ell}\right].$ (8)
-The first factor $e^{\ell(Z_{T})}/m(P_{\!T})$ is a probability weight under $P_{\!T}$ (it sums to 1).
-The second factor $e^{\Delta\ell}$ captures how the shift from $Z_{T}$ to $Z_{M}$ changes the score.
+The first term, $e^{\ell(Z_{T})}/m(P_{\!T})$, is a probability density under $P_{\!T}$ (it sums to 1).
+The second term, $e^{\Delta\ell}$ accounts for the change in the score when moving from $Z_{T}$ to $Z_{M}$.
 Since $|e^{x}-1|\leq|x|\cdot e^{|x|}$ for bounded $|x|\leq L_{\log}D$:
 $\left|\frac{m(P_{\!M})}{m(P_{\!T})}-1\right|\leq e^{L_{\log}D}\cdot\mathbb{E}_{\pi}[|\Delta\ell|]\leq e^{L_{\log}D}\cdot L_{\log}\cdot\mathbb{E}_{\pi}[\|Z_{M}-Z_{T}\|].$ (9)
-Using $|\log(1+x)|\leq|x|$ for $|x|<1$ (applicable when $W_{\!1}$ is small enough that the ratio stays near 1):
+Using $|\log(1+x)|\leq|x|$ for $|x|<1$ (applies when $W_{\!1}$ is small enough that the ratio stays near 1):
 $|B(P_{\!M})-B(P_{\!T})|=\left|\log\frac{m(P_{\!M})}{m(P_{\!T})}\right|\leq e^{L_{\log}D}\cdot L_{\log}\cdot W_{\!1}.$ (10)
-The $e^{L_{\log}D}$ prefactor is the price of the exponential: the competition term is more sensitive to distributional shift than the direct term because it involves $e^{\ell}$ rather than $\ell$.
-This is analogous to how compound interest amplifies small rate changes: a 1% shift in the base rate produces a much larger change in the compounded amount when the compounding period is long.
-Step 3: Combine. The total GMI change is bounded by the sum of the direct and competition terms:
+The $e^{L_{\log}D}$ term is the cost of the exponential: the competition term is more unstable to distributional shift than the direct term, because the former contains $e^{\ell}$ while the latter contains $\ell$.
+Step 3: Combine. The overall GMI change is at most the sum of the direct and competition terms:
 $|\Gamma(P_{\!M})-\Gamma(P_{\!T})|\leq\underbrace{L_{\log}\cdot W_{\!1}}_{\text{direct}}+\underbrace{e^{L_{\log}D}\cdot L_{\log}\cdot W_{\!1}}_{\text{competition}}=(1+e^{L_{\log}D})\cdot L_{\log}\cdot W_{\!1}.$ (11)
 The competition term dominates: it is $e^{L_{\log}D}$ times larger than the direct term.
-This is why the bound’s prefactor is large, and why the support-restricted corollary (which reduces $D$) is practically important.
+This is why the prefactor in the bound is so large and why the support-bounded corollary (which reduces $D$) is of practical value.
 Step 4: Average over $(C,Y)$.
 Everything above was for a fixed prompt $C$ and target $Y$.
-By Assumption 1 ($P_{\!M}(C,Y)=P_{\!T}(C,Y)$), the same prompts and targets appear under both laws, so we simply average over all $(C,Y)$ pairs to obtain the final bound.
+By Assumption 1 ($P_{\!M}(C,Y)=P_{\!T}(C,Y)$), the same prompts and targets appear under both laws, so we can simply average over all $(C,Y)$ pairs to obtain the final bound.
+When is the bound informative?
+The bound is a perturbation result: it is tightest when the mismatch $W_{\!1}$ is moderate relative to the decoder’s sensitivity.
+In the large-mismatch regime ($L_{\log}\cdot W_{\!1}\gg 1$), the prefactor $(1+e^{L_{\log}D})$ makes the right-hand side exceed the maximum possible GMI, and the bound becomes vacuous.
+In this regime, the bound correctly predicts that collapse occurs (the GMI ceiling is far below the matched-decoder rate) but cannot quantify how much information survives.
+Empirically, speech models operate in this regime ($L_{\log}\cdot W_{\!1}=19$$162$), which is why the bound is loose for speech but informative for vision ($L_{\log}\cdot W_{\!1}=13$$54$).
+The support-restricted corollary below partially addresses this by replacing the ambient diameter $D$ with the smaller effective diameter $D_{\text{eff}}$, reducing the prefactor from $\sim e^{150}$ to $\sim e^{7}$.
 A.3 Support-restricted bound
 Support-restricted bound.
-If $P_{\!M}(Z|c,y)$ and $P_{\!T}(Z|c,y)$ are both supported within a region $\mathcal{S}\subseteq\mathcal{Z}$ with $D_{\mathcal{S}}=\mathrm{diam}(\mathcal{S})\leq D$, then $D$ may be replaced by $D_{\mathcal{S}}$ in Theorem 2.
+If $P_{\!M}(Z|c,y)$ and $P_{\!T}(Z|c,y)$ are supported on some subset $\mathcal{S}\subseteq\mathcal{Z}$ such that $D_{\mathcal{S}}=\mathrm{diam}(\mathcal{S})\leq D$, then $D$ may be replaced by $D_{\mathcal{S}}$ in Theorem 2.
 Proof.
-If both marginals are supported in $\mathcal{S}$, any coupling $\pi$ has support on $\mathcal{S}\times\mathcal{S}$, so $\|Z_{M}-Z_{T}\|\leq D_{\mathcal{S}}$ for $\pi$-a.e.
+If both marginals are supported in $\mathcal{S}$, any coupling $\pi$ will also be supported on $\mathcal{S}\times\mathcal{S}$, so $\|Z_{M}-Z_{T}\|\leq D_{\mathcal{S}}$ for $\pi$-a.e.
 $(Z_{M},Z_{T})$.
-Step 2 above then bounds $|\Delta\ell|\leq L_{\log}\cdot D_{\mathcal{S}}$, giving the tighter prefactor.
+Step 2 above then gives that $|\Delta\ell|\leq L_{\log}\cdot D_{\mathcal{S}}$, which is the improved constant factor.
 Why this matters.
-The ambient diameter $D$ makes the prefactor $(1+e^{L_{\log}D})$ astronomically large ($L_{\log}\cdot D$ ranges from 10 to 150).
-But modal and text representations often occupy a much smaller effective region.
-We estimate the effective support diameter $D_{\text{eff}}$ via the participation ratio of the pooled representation covariance (§5.5), yielding $L_{\log}\cdot D_{\text{eff}}\in[3,7]$ and a moderate prefactor ($\sim$20–1100).
-A.4 Probe-decoder asymmetry
+The ambient diameter $D$ makes the prefactor $(1+e^{L_{\log}D})$ extremely large ($L_{\log}\cdot D$ ranges from 10 to 150), while representations in modal and text space frequently lie in a far smaller effective volume.
+We estimate the effective support diameter $D_{\text{eff}}$ from the participation ratio of the pooled representation covariance (§5.6), yielding $L_{\log}\cdot D_{\text{eff}}\in[3,7]$ and a moderate prefactor ($\sim$20–1100).
+A.4 Per-type prediction
+The theory predicts the overall trend: text-correlated information types (lexical) should strengthen or hold steady through the LLM, while non-text types (speaker, emotion, acoustic) should weaken.
+This qualitative prediction holds across all five models.
+A stronger prediction would be that the degree of weakening for each information type matches its share of modality-specific variance.
+This does not hold at the level of individual types (Spearman $\rho=-0.40$, $p=0.60$ across the four speech information types in Ultravox).
+The likely cause is the content-dependent $L_{\log}$: the Lipschitz constant varies $2.5\times$ across audio types ($L_{\log}=14.58$ for speech vs.
+$5.92$ for environmental sounds in Ultravox), acting as an additional driver of degradation independent of geometric alignment.
+A per-type bound would need to incorporate this content-dependence, replacing the scalar $L_{\log}$ with a type-conditional $L_{\log}(\tau)$.
+A.5 Probe-decoder asymmetry
 Theorem 3 (Probe penalty).
@@
 Proof.
-Identical to Step 1 of the Theorem 2 proof, replacing $\ell_{\psi}$ with $\log h$ and $L_{\log}$ with $L_{h}$.
-Crucially, there is no competition term for probes (no random-coding argument), so the exponential prefactor does not appear.
+Identical to Step 1 of the Theorem 2 proof, replacing $\ell$ with $\log h$ and $L_{\log}$ with $L_{h}$.
+Also there is no competition term for probes (no random-coding argument), so the exponential prefactor does not appear.
 The key asymmetry.
-A linear probe is a simple instrument: one matrix multiplication followed by a softmax.
-Its sensitivity to distributional shift ($L_{h}$) is just the size of that matrix’s weights.
+A linear probe is a simple model: one matrix multiplication followed by a softmax.
+Its sensitivity to distributional shift ($L_{h}$) is therefore simply the norm of the weights of that matrix.
 The LLM decoder is a deep transformer with billions of parameters, and its sensitivity ($L_{\log}$) reflects that depth.
-Empirically, $L_{\log}/L_{h}\approx 30\times$5.5).
-Now consider the same distributional shift (modal representations replacing text representations).
+Empirically, $L_{\log}/L_{h}\approx 30\times$5.6).
+Now consider the same distributional shift with modal representations replacing text representations.
 The probe’s bound says degradation $\leq L_{h}\cdot W_{\!1}$ (small).
 The decoder’s bound says degradation $\leq(1+e^{L_{\log}D})\cdot L_{\log}\cdot W_{\!1}$ (large).
-Same shift, dramatically different impact.
+Same shift, different impact.
 This is the formal basis for the “present but inaccessible” phenomenon: a probe can still read the information because it is a simple, low-sensitivity instrument, but the decoder’s complex scoring rule is far more sensitive to the distributional mismatch and cannot use it.
@@
 $L_{\log}$ estimated via the 2-nearest-neighbor gradient norm method over n=1,000n=1{,}000 samples.
-For each sample $z_{i}$, we compute $\|\nabla_{z}\log q_{\psi}(y_{i}|z_{i},c_{i})\|$ via autograd.
+For each sample $z_{i}$, we compute $\|\nabla_{z}\log q(y_{i}|z_{i},c_{i})\|$ via autograd.
 The 95th percentile of the resulting distribution serves as the $L_{\log}$ estimate.
 B.3 Extraction prompts
-During representation extraction, each model receives a minimal, task-neutral prompt to avoid biasing hidden states toward any particular downstream task.
-Representations are extracted from intermediate hook points, not from the generated output.
+For the representation extraction experiments, we provide each model with a minimal, task-agnostic input to prevent the hidden states from being biased toward a specific downstream task.
+The representations are always extracted from the intermediate hook locations, not from the output.
 Speech models.
@@
 The emotion LoRA (§6.1) uses standard causal LM fine-tuning on Ultravox with a forced-choice emotion detection objective on CREMA-D.
-LoRA configuration. Rank r=16r{=}16, $\alpha{=}32$, dropout $0.05$, target modules: q_proj, k_proj, v_proj, o_proj (same configuration as the ALME LoRA used as the negative example).
+LoRA configuration. Rank r=16r{=}16, $\alpha{=}32$, dropout $0.05$, target modules: q_proj, k_proj, v_proj, o_proj.
 Training.
@@
   • Assistant: {"answer": "<emotion>"} (training target)
-Loss is computed only on the assistant response tokens; all input/prompt tokens are masked with label 100-100.
+We calculate the loss only on the assistant response tokens, while the input/prompt tokens are masked (with label 100-100).
 Data split. CREMA-D (7,442 clips, 91 speakers, 6 emotions). 80/20 stratified split yields 1,002 held-out evaluation samples (167 per emotion class), saved deterministically for reproducibility.
-Baseline evaluation. Base Ultravox (no LoRA) achieves $17.3\%$ generation accuracy on the held-out split (near chance for 6 classes), defaulting to “neutral” for most inputs.
+Baseline evaluation. Base Ultravox (no LoRA) achieves $17.3\%$ task accuracy on the held-out split (near chance for 6 classes), defaulting to “neutral” for most inputs.
 B.5 Full probe accuracy tables
 Table 5 reports the full probe accuracies (mean $\pm$ std over 5 seeds) at all four hook points for every model–dataset–information-type combination.
-The key pattern to observe: probe accuracy for non-text information types remains well above chance even at the LLM’s final layer, confirming that the information survives the decoder’s processing.
+The main finding is that probe accuracy for non-text information types remains well above chance even at the LLM’s final layer, confirming that the information survives the decoder’s processing.
 It is the decoder’s use of this information, not its presence, that fails.
@@
 Values above 100% indicate the LLM amplifies the signal; values below 100% indicate degradation.
-The key pattern: non-aligned encoders show asymmetric behavior (lexical content is amplified while speaker identity collapses), whereas text-aligned encoders retain or slightly improve all information types, but only because the encoder has already discarded non-textual information.
+The pattern to notice is that non-text aligned encoders exhibit an asymmetric pattern (lexical content is amplified while speaker identity degrades), while text aligned encoders preserve or improve all the information types, simply because the encoder has already filtered out non-text information.
 Table 6: Information retention (adapter $\to$ LLM-final, %).
@@
 Restricting to audio tokens does not help, and in several cases slightly hurts.
-This is because the LLM’s self-attention layers redistribute information across all sequence positions as they process the input.
-By layer 16, audio-derived information has spread to text prompt tokens as well.
-Pooling over the full sequence therefore captures more of the available signal, not less.
-This validates our use of full-sequence pooling throughout the paper: it is not a conservative approximation that dilutes the audio signal, but rather the appropriate strategy given how transformers distribute information.
+This is consistent with the LLM mixing information across sequence positions during processing, so that the audio-derived signal is available at non-audio positions by layer 16.
 Dataset Label All-token Audio-only $\Delta$
@@
 […]
-Table 9: Probe accuracy with all-token vs.
-audio-only pooling at LLM layers (Ultravox).
-All deltas are within $\pm$2.1%, most under 1%.
-Audio-only pooling does not improve probe accuracy.
-The LLM’s self-attention layers mix information across all sequence positions: by layer 16, audio-derived information has been redistributed to text prompt tokens as well.
-Restricting pooling to audio positions therefore discards information that the LLM has already spread across the full sequence, slightly reducing the available signal.
-This validates our use of full-sequence pooling throughout the paper.
-Appendix C Additional figures
-Figure 1: Probe accuracy trajectories for all five models.
-Rows 1–2: speech models (Ultravox, Qwen2-Audio) across LibriSpeech, CREMA-D, and ESC-50.
-Row 3: vision models (LLaVA, Prismatic-DINOv2, Prismatic-SigLIP) on COCO.
+Table 9: Probe accuracy with all-token vs. audio-only pooling at LLM layers (Ultravox). All deltas are within $\pm$2.1%, most under 1%. Audio-only pooling does not improve probe accuracy.
+Figure 1: Linear probe accuracy at four extraction points (encoder, adapter, LLM layer 16, LLM final) for all five models.
+For speech models (rows 1–2), lexical accuracy increases through the LLM while speaker identity degrades, illustrating the information accessibility gap.
+For vision models (row 3), LLaVA and Prismatic-S (text-aligned encoders) show improvement across all information types, while Prismatic-D (non-aligned) shows flat or declining non-textual accuracy.
+Datasets: LibriSpeech, CREMA-D, ESC-50 (speech); COCO (vision).
 Figure 2: Mode alignment profiles for all five models.
-Blue: alignment score $\tilde{\alpha}(u_{k})$; orange: eigenvalue spectrum $\lambda_{k}$ (log scale).
-The dominant eigenmodes are modality-specific ($\tilde{\alpha}\approx 0$) for all models with non-text-aligned encoders; for Prismatic-S (SigLIP, text-aligned), even Mode 0 is text-aligned ($\tilde{\alpha}=0.83$).
+Each panel shows the top modes (eigenvectors of the adapter-output covariance $\Sigma_{M}$, sorted by decreasing eigenvalue; 100 for speech models, 200 for vision).
+Blue (left axis): alignment score $\rho_{k}$, the ratio of text variance to modal variance along each mode (§5.3).
+Orange (right axis): eigenvalue $\lambda_{k}$ (log scale), the amount of adapter variance carried by each mode.
+Annotations show the percentage of total variance in Mode 0 and the count of modality-specific modes ($\rho_{k}<0.5$).
+For non-aligned encoders, the highest-variance modes have $\rho_{k}\approx 0$: the adapter’s dominant variation is along directions the decoder has never seen.
+For Prismatic-S (SigLIP, text-aligned), even Mode 0 is text-aligned ($\rho_{0}=0.83$).
 The Prismatic comparison (Table 10) isolates encoder text-alignment as the causal variable.
@@
 For each model, we decompose the adapter output covariance into eigenmodes and measure their text alignment.
-The dominant eigenmode is modality-specific ($\tilde{\alpha}\ll 1$) for all non-text-aligned encoders, while for SigLIP even the dominant mode is text-aligned ($\tilde{\alpha}=0.83$).
-Model Mode 0 $\tilde{\alpha}$ Mean $\tilde{\alpha}$ (top 10) MS modes MS var.
+The dominant eigenmode is modality-specific ($\rho_{k}\ll 1$) for all non-text-aligned encoders, while for SigLIP even the dominant mode is text-aligned ($\rho_{k}=0.83$).
+Model Mode 0 $\rho_{k}$ Mean $\rho_{k}$ (top 10) MS modes MS var.
 Ultravox 0.001 1.76 11/100 63.6%
@@
 Table 11: Mode alignment summary.
-MS modes = modes with $\tilde{\alpha}<0.5$ (modality-specific).
+MS modes = modes with $\rho_{k}<0.5$ (modality-specific).
 MS var.
 = fraction of top-100 eigenmode variance in MS modes.
-Mode 0 is modality-specific ($\tilde{\alpha}\ll 1$) for all non-text-aligned encoders; for SigLIP (text-aligned), even Mode 0 is text-aligned ($\tilde{\alpha}=0.83$).
+Mode 0 is modality-specific ($\rho_{k}\ll 1$) for all non-text-aligned encoders; for SigLIP (text-aligned), even Mode 0 is text-aligned ($\rho_{k}=0.83$).
 Extending to all 4096 modes strengthens the pattern: tail modes are overwhelmingly text-aligned with negligible individual variance, so the MS/TA split is driven by the high-variance modes reported here.
-Appendix D Additional limitations
+Appendix C Additional limitations
 Scope. We test speech and vision across five models. The framework applies to any multimodal LLM with a text-trained decoder, but video/3D remain untested.
@@
 Generality of the GMI constraint.
-The mechanism described in the main text is not architectural: any multimodal LLM whose decoder was trained under a text-dominated objective has a text-shaped scoring rule, regardless of whether the adapter is a linear projection, MLP, Q-Former, perceiver resampler, or discrete codebook.
-The GMI constraint applies to multimodal modelling in general: unless the training objective explicitly targets each modality, the decoder remains indifferent to non-text directions.
+The mechanistic explanation is not architectural: the text-shaped scoring rule of the multimodal LLM decoder is conditioned only on the text-dominating training objective, and holds regardless of the adapter implementation (linear projection, MLP, Q-Former, perceiver resampler, discrete codebook, etc.).
+The GMI constraint is more broadly relevant for multimodal modelling: in the absence of a training objective that explicitly seeks out each modality, the decoder will remain indifferent to non-text directions.